### v1

# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

Abstract.
The validity of functional inequalities on Finsler metric measure manifolds is based on three non-Riemannian quantities, namely, the reversibility, flag curvature and $S$-curvature induced by the measure. Under mild assumptions on the reversibility and flag curvature, it turned out that famous functional inequalities – as Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality – usually hold on forward complete Finsler manifolds with non-positive $S$-curvature, cf. Huang, Kristály and Zhao [Trans. Amer. Math. Soc., 2020]. In this paper however we prove that – under similar assumptions on the reversibility and flag curvature as before – the aforementioned functional inequalities fail whenever the $S$-curvature is positive. Accordingly, our results clearly reveal the deep dependence of functional inequalities on the $S$-curvature. As a consequence of these results, we establish surprising analytic aspects of Finsler manifolds: if the flag curvature is non-positive, the Ricci curvature is bounded from below and the $S$-curvature is positive, then the reversibility turns out to be infinite. Examples are presented on general Funk metric spaces, where the $S$-curvature plays again a decisive role.
Key words and phrases:
Finsler manifold; $S$-curvature; flag curvature; Ricci curvature; reversibility; functional inequalities
2010 Mathematics Subject Classification:
Primary 26D10, Secondary 53C60, 53C23
A. Kristály is supported by the Excellence Researcher Program ÓE-KP-2-2022 of Óbuda University, Hungary. B. Li is supported by Natural Science Foundation of Zhejiang Province (No. LY23A010012), Natural Science Foundation of Ningbo (2021J110) and K.C. Wong Magna Fund in Ningbo University. W. Zhao is supported by Natural Science Foundation of Shanghai (No. 21ZR1418300) and Science and Technology Project of Xinjiang Production and Construction Corps (No. 2023CB008-12).
1. Introduction
As the title of S.S. Chern’s famous 1996 paper [7] shows, his credo was that ‘Finsler geometry is just Riemannian geometry without the quadratic restriction’. In addition, Chern concluded his paper by claiming that ”(…) almost all the results of Riemannian geometry can be developed in the Finsler setting. It is remarkable that one needs only some conceptual adjustment, no essential new ideas being necessary.” To a certain point, Chern’s position cannot be discredited, since most of the well-known results from Riemannian geometry – by suitable adaptations – have their natural Finslerian counterparts, e.g., Hopf–Rinow, Cartan–Hadamard, Gauss–Bonnet and Bonnet–Myers theorems as well as Rauch, Bishop–Gromov and Bishop–Günther comparison principles; see Auslander [1], Bao, Chern and Shen [2], Rund [25] and Shen [27]. However, the development of Finsler geometry over the last three decades clearly shows that fundamentally new ideas are needed to solve deep problems, revealing in this way genuine differences between Riemannian and Finsler geometries. In support of this thesis, we recall some open questions and spectacular results from Finsler geometry, the solutions to which are far from being tied to the Riemannian world:
  • Classification of Finsler manifolds with constant flag curvature. Unlike the Hopf classification in Riemannian geometry, no full characterization is available for Finsler manifolds having constant flag curvature; in fact, it seems that different subclasses of Finsler manifolds – as Minkowski, Berwald, Landsberg, projectively flat and Randers spaces – play a crucial role in such a characterization, see e.g., Bao, Robles and Shen [3], Li [19], and Shen [28, 29, 31].
  • Busemann NPC spaces. Busemann non-positively curved (shortly, NPC) spaces on Riemannian manifolds are characterized by the non-positivity of the sectional curvature, see Busemann [4] and Jost [13]. It was shown however by Kelly and Straus [15] that any strictly convex domain $D\subset\mathbb{R}^{2}$ endowed with the Hilbert metric (which is a projectively flat Finsler manifold with constant negative curvature) is a Busemann NPC space if and only if the boundary of $D$ is an ellipse and the Hilbert geometry is hyperbolic. Ivanov and Lytchak [12] proved that a Finsler manifold is a Busemann NPC space if and only if it is a Berwald manifold of non-positive flag curvature.
  • Sobolev spaces over Finsler manifolds. Contrary to the Riemannian setting, the Sobolev spaces over noncompact, irreversible Finsler manifolds should not have even a vector space structure, see Kristály and Rudas [18]. This construction was carried out over the Funk metric defined on the $n$-dimensional Euclidean unit ball $\mathbb{B}^{n}$.
In addition to the above list, the present paper provides another new aspect of non-Riemannian Finsler manifolds: we show that famous functional inequalities – which naturally hold on Riemannian manifolds – fail on a large class of Finsler manifolds.
Let $(M,F,\operatorname{\mathfrak{m}})$ be a Finsler metric measure manifold ($\operatorname{\mathsf{FMMM}}$, for short), i.e., a Finsler manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$, see Section 2 for details. Huang, Kristály and Zhao [10] already pointed out that three non-Riemannian quantities strongly influence the validity of functional inequalities on $(M,F,\operatorname{\mathfrak{m}})$, namely,
  • reversibility;
  • $S$-curvature induced by the measure $\operatorname{\mathfrak{m}}$;
  • flag curvature.
The reversibility of a Finsler manifold $(M,F)$, introduced by Rademacher [24], is defined as
$\lambda_{F}(M):=\sup_{x\in M}\lambda_{F}(x),\quad\text{ where }\quad\lambda_{F}(x):=\sup_{y\in T_{x}M\backslash\{0\}}\frac{F(x,-y)}{F(x,y)}.$
It is easy to see that $\lambda_{F}(M)\geq 1$ with equality if and only if $F$ is reversible (i.e., symmetric). Obviously, all Riemannian metrics are reversible, but there are infinitely many irreversible Finsler metrics. Typical examples with infinite reversibility are Funk metric spaces, which served to prove that Sobolev spaces over such Finsler manifolds are not linear structures, see Farkas, Kristály and Varga [9] and Kristály and Rudas [18]. Even if the Finsler metric is explicitly given, it is usually a hard task to determine whether the reversibility is finite or not.
Unlike in the Riemannian case, there is no a canonical measure on Finsler manifolds. Thus, various measures can be introduced whose behavior may be genuinely different. Every smooth positive measure $\operatorname{\mathfrak{m}}$ on a Finsler manifold $(M,F)$ induces two non-Riemannian quantities, which are the so-called distortion $\tau$ and $S$-curvature $\mathbf{S}$, respectively, see Shen [27]. More precisely, set $\operatorname{d\mathfrak{m}}:=\sigma(x){\operatorname{d}}x^{1}\wedge\cdots\wedge{\operatorname{d}}x^{n}$ in some local coordinate system $(x^{i})$. Thus, for any $y\in T_{x}M\backslash\{0\}$, define
$\tau(x,y):=\ln\frac{\sqrt{\det g_{ij}(x,y)}}{\sigma(x)},\qquad\mathbf{S}(x,y):=\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\tau(\gamma_{y}(t),{\gamma}^{\prime}_{y}(t)),$
where $g_{y}=(g_{ij}(x,y))$ is the fundamental tensor of $F$ and $t\mapsto\gamma_{y}(t)$ is the geodesic starting from $x$ with initial speed $\gamma^{\prime}_{y}(0)=y$. Since ${\bf S}$ is 1-positive homogeneous, we say that $\mathbf{S}\geq a$ for some constant $a\in\mathbb{R}$ if $\mathbf{S}(x,y)\geq aF(x,y)$ for all $(x,y)\in TM\backslash\{0\}.$
For each $y\in T_{x}M$ and any plane $\Pi=\text{span}\{y,v\}\subset T_{x}M$, the flag curvature is defined by
$\mathbf{K}(y;\Pi):=\mathbf{K}(y;v):=\frac{g_{y}(R_{y}(v),v)}{F^{2}(x,y)\,g_{y}(v,v)-g^{2}_{y}(v,y)^{2}},$
where $R_{y}$ is the Riemannian curvature of $F.$ Clearly, the flag curvature naturally defines the Ricci curvature ${\bf Ric}.$ As expected, on any Riemannian manifold endowed with its canonical measure the $S$-curvature vanishes, while the flag curvature reduces to the sectional curvature.
Before stating our main results, we recall the main achievements from Huang, Kristály and Zhao [10], which represent the starting points of our present study; in fact, they roughly state that if
${\bf K}\leq 0\ \ {\rm and}\ \ {\bf S}\leq 0,$ (1.1)
then the well-known functional inequalities hold with sharp (Euclidean) constants, including the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality, respectively.
To be more explicit, let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional Cartan–Hadamard $\operatorname{\mathsf{FMMM}}$ (i.e., simply connected, forward complete $\operatorname{\mathsf{FMMM}}$ with ${\bf K}\leq 0$) and ${\bf S}\leq 0$. If $r(x)=d_{F}(o,x)$ is the distance function from a point $o\in M$, and $F^{*}$ stands for the co-metric of $F$, then the following functional inequalities hold (for their purely Riemannian versions, see Kristály [16]):
  • Hardy inequality; see [10, Theorem 1.4 for $p=2$] and [36]. If $p\in(1,n),$ then
    ${\displaystyle\int_{M}\max\left\{{F}^{*p}({\pm\rm d}u)\right\}\operatorname{d\mathfrak{m}}}\geq\left(\frac{n-p}{p}\right)^{p}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}},\ \forall u\in C^{\infty}_{0}(M);$ (1.2)
    moreover, if $F$ is reversible, then the constant $\left(\frac{n-p}{p}\right)^{p}$ is sharp and never achieved.
  • Heisenberg–Pauli–Weyl uncertainty principle; see [10, Theorem 1.1]. If $n\geq 2,$ then
    $\ \ \ \ \ {\displaystyle\left(\int_{M}\max\left\{{F}^{*2}({\pm\rm d}u)\right\}\operatorname{d\mathfrak{m}}\right)}\left(\int_{M}{r^{2}u^{2}}\operatorname{d\mathfrak{m}}\right)\geq\frac{n^{2}}{4}{\displaystyle\left(\int_{M}{u^{2}}\operatorname{d\mathfrak{m}}\right)^{2}},\ \forall u\in C^{\infty}_{0}(M);$ (1.3)
    in addition, if $F$ is reversible, then $\frac{n^{2}}{4}$ is sharp and is achieved by a non-trivial function if and only if $\operatorname{\mathfrak{m}}$ is the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{BH}$, ${\bf K}=0$ and ${\bf S}={\bf S}_{BH}=0$.
  • Caffarelli–Kohn–Nirenberg interpolation inequality; see [10, Theorem 1.1]. If $0<q<2<p$ and $2<n<2(p-q)/(p-2)$, then for every $u\in C^{\infty}_{0}(M)$ one has
    ${\displaystyle\left(\int_{M}\max\left\{{F}^{*2}({\pm\rm d}u)\right\}\operatorname{d\mathfrak{m}}\right)}\left(\int_{M}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)\geq\frac{(n-q)^{2}}{p^{2}}{\displaystyle\left(\int_{M}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}};$ (1.4)
    furthermore, if $F$ is reversible, then $\frac{(n-q)^{2}}{p^{2}}$ is sharp and is achieved as in the case (1.3).
On the other hand, if $(M,F,\operatorname{\mathfrak{m}})$ is an $n$-dimensional reversible $\operatorname{\mathsf{FMMM}}$, and the ’complementary’ case to (1.1) holds, i.e.,
${\bf Ric}\geq 0\ \ {\rm and}\ \ {\bf S}\geq 0,$
together with either the Heisenberg–Pauli–Weyl uncertainty principle (1.3) or the Caffarelli–Kohn–Nirenberg interpolation inequality (1.4), then we have the rigidity that $\operatorname{\mathfrak{m}}$ is the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{BH}$ (up to a multiplicative constant), ${\bf K}=0$ and ${\bf S}={\bf S}_{BH}=0$, see [10, Theorem 1.3].
The above set of results contain in certain sense the genesis of our first main result – which reads as follows – and roughly shows that the validity/failure of functional inequalities on Finsler manifolds is decided by the sign of the $S$-curvature:
Theorem 1.1.
Let $n\geq 2$ be an integer, $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$, and $r=d_{F}(o,\cdot)$ for some $o\in M$. Suppose that
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h},$ (1.5)
where $k,h$ are two constants with $h>k\geq 0$. Then the following statements hold::::
  1. (i)
    the first eigenvalue of the $p$-Laplacian vanishes on $(M,F,\operatorname{\mathfrak{m}})$ for every $p>1$, i.e.,
    $\inf_{u\in C^{\infty}_{0}(M)}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}{|u|^{p}}\operatorname{d\mathfrak{m}}}=0;$
  2. (ii)
    the $L^{p}$-Hardy inequality fails on $(M,F,\operatorname{\mathfrak{m}})$ for every $p>1$, i.e.,
    $\inf_{u\in C^{\infty}_{0}(M)}\frac{\displaystyle\int_{M}{F}^{*p}({\rm d}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0;$
  3. (iii)
    the Heisenberg–Pauli–Weyl principle fails on $(M,F,\operatorname{\mathfrak{m}})$, i.e.,
    $\inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}r^{2}u^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0;$
  4. (iv)
    the Caffarelli–Kohn–Nirenberg interpolation inequality fails on $(M,F,\operatorname{\mathfrak{m}})$, i.e., if $0<q<2<p$ and $2<n<2(p-q)/(p-2)$, then
    $\inf_{u\in C^{\infty}_{0}({M})\backslash\{0\}}\frac{\left(\displaystyle\int_{{M}}F^{*2}({\rm d}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{{M}}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{{M}}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$
Intuitively, the $S$-curvature describes the change of measure along geodesics, see Shen [30]. As far as we know, only a few results are available in the literature concerning the impact of $S$-curvature on global properties of Finsler manifolds. In the light of the results from [10], Theorem 1.1 provides a concluding answer about the validity of functional inequalities in Finsler manifolds, which is captured by the behavior of the $S$-curvature. In particular, all the statements except (i) in Theorem 1.1 will be invalid if the condition $h>0$ is relaxed to $h\geq 0$, showing in this way the optimality of our results. Indeed, every reversible Minkowski space endowed with the Busemann–Hausdorff measure is a Cartan–Hadamard manifold with $\mathbf{K}=0$ and $\mathbf{S}=0$ (i.e., $h=k=0$); thus the (sharp) inequalities from (1.2)-(1.4) all are valid, which contradict the conclusions from Theorem 1.1. We also notice that there are large classes of non-Riemannian reversible Minkowski spaces; see Example 6.1.
The proof of Theorem 1.1 is obtained by considering polar coordinates on $M$ with a suitable disintegration of the measure $\operatorname{\mathfrak{m}}$, combined with the key assumption (1.5) with $h>k\geq 0$, comparison principles and suitably chosen test functions (with appropriate approximation arguments). For the proof of (i)-(iii) the Guassian bubble $u_{\alpha}=-e^{-\alpha r}$ is used, while for (iv) a Talentian bubble is considered. The Gaussian bubble has been used in Kristály [17] and Huang, Kristály and Zhao [10] to prove (i) and (ii), respectively, in the special case of the Funk metric on the Euclidean unit ball $\mathbb{B}^{n}$. In general, according to Li [19] and Shen [31, 26], for every non-empty bounded strongly convex domain $\Omega\subset\mathbb{R}^{n}$ (containing the origin $\mathbf{0}$) with smooth boundary, there always exists a Finsler metric $F$, which is a Funk metric, and $(\Omega,F)$ is a Cartan–Hadamard manifold with $\mathbf{K}=-1/4$, see Section 6. Such a pair $(\Omega,F)$ is called a Funk metric space, and it is the model space where our Theorem 1.1 can be applied. In the particular case when $\Omega=\mathbb{B}^{n}$ and $\operatorname{\mathfrak{m}}=\operatorname{\mathfrak{m}}_{BH}$, the Funk metric on $\mathbb{B}^{n}$ is
$F(x,y)=\frac{\sqrt{|y|^{2}-|x|^{2}|y|^{2}+\langle x,y\rangle^{2}}+\langle x,y\rangle}{1-|x|^{2}},\quad x\in\mathbb{B}^{n},\ y\in T_{x}\mathbb{B}^{n}=\mathbb{R}^{n},$ (1.6)
see Shen [31, 26, 27], for which ${\bf Ric}=-(n-1)/4$ and ${\bf S}_{BH}=(n+1)/2,$ thus $h=\frac{n+1}{2(n-1)}>\frac{1}{2}=k>0.$ However, the symmetrization of $F$ is $\tilde{F}(x,y)=(F(x,y)+F(x,-y))/2,$ which is precisely the Riemannian Hilbert/Klein metric on $\mathbb{B}^{n}$ with ${\bf K}=-1$ (so ${\bf Ric}=-(n-1)$) and ${\bf S}=0$; thus the inequalities (1.2)-(1.4) apply instead of Theorem 1.1.
In the sequel, we investigate the influence of $S$-curvature on analytic properties of Finsler manifolds; more precisely, it turns out that the positivity of $S$-curvature has a strong implication on the reversibility.
Theorem 1.2.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete noncompact $\operatorname{\mathsf{FMMM}}$. Assume that
$\mathbf{K}\leq 0,\quad\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h},$ (1.7)
where $k,h$ are two constants satisfying $h>0$ and $h\geq k\geq 0$. Then $\lambda_{F}(M)=+\infty$.
Obviously, every Funk metric space endowed with the Busemann-Hausdorff measure verifies the assumptions of Theorem 1.2. The additional condition $\mathbf{K}\leq 0$ is mild. In fact, if the flag curvature is bounded from below by a positive number, then the Bonnet–Myers theorem implies the compactness of the manifold, see [2, Theorem 7.7.1], in which case the reversibility is finite.
We recall that another important notion in Finsler geometry is the weighted Ricci curvature $\mathbf{Ric}_{N}$ for $N\in[n,\infty]$. Although $\mathbf{Ric}_{N}$ is a combination of $\mathbf{Ric}$ and $\mathbf{S}$, we show that one cannot replace the separated conditions from (1.5) (or (1.7)) by $\mathbf{Ric}_{N}\geq-K^{2}$ for any $K\geq 0$. In fact, every reversible Minkowski space equipped with the Busemann–Hausdorff measure satisfies $\mathbf{Ric}_{N}=0$, but neither Theorem 1.1 nor Theorem 1.2 holds on such spaces; for further examples, see subsection 6.2. Moreover, it is not appropriate either to assume $\mathbf{Ric}_{N}\geq K>0,$ since in this case the manifold is compact and consequently, its reversibility is finite.
The paper is organized as follows. Section 2 is devoted to preliminaries on Finsler geometry together with some results about volume comparison. In Section 3, we provide some immediate properties of Sobolev spaces over irreversible Finsler manifolds. Here, we also establish an approximation result which will be useful in our main results. The proof of our main Theorem 1.1 is given in Section 4, while the proof of Theorem 1.2 is presented in Section 5. Section 6 is devoted to the study of Funk metric spaces, by providing examples and counterexamples supporting the sharpness of our main results.
2. Preliminaries
2.1. Elements from Finsler geometry
In this section we recall some definitions and properties from Finsler geometry; for details see e.g. Bao, Chern and Shen [2] and Shen [26, 27].
2.1.1. Finsler manifolds
A Minkowski norm on an $n$($\geq 2$)-dimensional vector space $V$ is a $C^{\infty}$-function $\phi:V\setminus\{0\}\rightarrow[0,+\infty)$ satisfying:
  1. (i)
    $\phi(y)\geq 0$ with equality if and only if $y=0$;
  2. (ii)
    $\phi$ is positively homogeneous of degree one, i.e., $\phi(\alpha y)=\alpha\phi(y)$ for all $\alpha>0$;
  3. (iii)
    $\phi(y)$ is strongly convex, i.e., the matrix $g_{ij}(y):=\frac{1}{2}\frac{\partial^{2}\phi^{2}}{\partial y^{i}\partial y^{j}}(y)$ is positive definite;
having these properties, the pair $(V,\phi)$ is called a Minkowski space.
Let $M$ be an $n$-dimensional (connected) smooth manifold and $TM=\bigcup_{x\in M}T_{x}M$ be its tangent bundle. A Finsler metric $F=F(x,y)$ on $M$ is a $C^{\infty}$ function on $TM\setminus\{0\}$ such that $F|_{T_{x}M}$ is a Minkowski norm on $T_{x}M$ for each $x\in M$, in which case the pair ($M$, $F$) is called a Finsler manifold.
The fundamental tensor is defined by
$g_{y}:=\big{(}g_{ij}(x,y)\big{)}=\left(\frac{1}{2}\frac{\partial^{2}F^{2}}{\partial y^{i}\partial y^{j}}(x,y)\right),\quad\forall\,(x,y)\in TM\backslash\{0\},$ (2.1)
which induces a Riemannian metric on $T_{x}M\backslash\{0\}$. The Finsler metric $F$ is (induced by) a Riemannian metric if and only if $g_{y}$ is independent of $y$, i.e., $g_{ij}(x,y)=g_{ij}(x)$. In the same way, $F$ is a Minkowski metric if $g_{y}$ is independent of $x$, i.e., $g_{ij}(x,y)=g_{ij}(y)$.
The reversibility of a set $U\subset M$, cf. Rademacher [24], is defined by
$\lambda_{F}(U):=\sup_{x\in U}\lambda_{F}(x),\quad\text{where}\quad\lambda_{F}(x)=\sup_{y\in T_{x}M\setminus\{0\}}\frac{F(x,-y)}{F(x,y)}.$ (2.2)
Thus $\lambda_{F}(M)\geq 1,$ with equality if and only if $F$ is reversible, i.e., $F(x,y)=F(x,-y)$ for every $(x,y)\in TM$. Moreover, $x\mapsto\lambda_{F}(x)$ is continuous on $M$. In particular, $\lambda_{F}(M)$ is finite when $M$ is compact or $F$ is a Minkowski metric.
The Legendre transformation $\mathfrak{L}:TM\rightarrow T^{*}M$ is defined by
$\mathfrak{L}(x,y):=\begin{cases}g_{y}(y,\cdot),&\text{ if }y\in T_{x}M\backslash\{0\},\\ 0,&\text{ if }y=0\in T_{x}M,\end{cases}$
where $g_{y}(y,\cdot):=g_{ij}(x,y)y^{i}{\operatorname{d}}x^{j}$. Given a smooth function $u:M\rightarrow\mathbb{R}$, the gradient of $u$ is
$\nabla u(x)=\mathfrak{L}^{-1}({\operatorname{d}}u).$
In particular, for any smooth vector filed $X$, one has $\langle X,{\operatorname{d}}u\rangle=X(u)=g_{\nabla u}(\nabla u,X),$ where $\langle y,\xi\rangle:=\xi(y)$ denotes the canonical pairing between $T_{x}M$ and $T^{*}_{x}M$.
The co-metric $F^{*}$ of $F$ on $M$ is defined by
$F^{*}(x,\xi):=\sup_{y\in T_{x}M\setminus\{0\}}\frac{\xi(y)}{F(x,y)},\quad\forall\xi\in T^{*}_{x}M,$ (2.3)
which is a Finsler metric on the cotangent bundle $T^{*}M$. It is easy to see that
$\langle y,\xi\rangle\leq F(x,y)F^{*}(x,\xi),\quad\text{ for any }y\in T_{x}M,\ \xi\in T^{*}_{x}M,$ (2.4)
with equality if and only if $\xi=\alpha\mathfrak{L}(x,y)$ for some $\alpha>0$. In addition, $F(x,y)=F^{*}(\mathfrak{L}(x,y))$ for any $y\in T_{x}M$. Furthermore, the reversibilities of $F$ and $F^{*}$ coincide, see Huang, Kristály and Zhao [10, Lemma 2.1], i.e.,
$\lambda_{F}(x)=\lambda_{F^{*}}(x),\quad\forall\,x\in M.$ (2.5)
A smooth curve $t\mapsto\gamma(t)$ in $(M,F)$ is called a geodesic if it satisfies
$\frac{{\operatorname{d}}^{2}\gamma^{i}}{{\operatorname{d}}t^{2}}+2G^{i}\left(\gamma,\frac{{\operatorname{d}}\gamma}{{\operatorname{d}}t}\right)=0,$ (2.6)
where $G^{i}=G^{i}(x,y)$ are the geodesic coefficients (or spray coefficients) given by
$G^{i}=\frac{1}{4}g^{ij}\left\{\frac{\partial^{2}F^{2}}{\partial x^{l}\partial y^{j}}y^{l}-\frac{\partial F^{2}}{\partial x^{j}}\right\},$ (2.7)
$(g^{ij})$ being the inverse matrix of $(g_{ij})$. In the sequel, we always use $\gamma_{y}(t)$, $t\geq 0,$ to denote the geodesic with initial velocity $y$.
For each $y\in T_{x}M$ and any plane $\Pi=\text{span}\{y,v\}\subset T_{x}M$, the flag curvature is defined by
$\mathbf{K}(y;\Pi):=\mathbf{K}(y;v):=\frac{g_{im}(x,y)\,R^{i}_{\ k}(x,y)\,v^{k}v^{m}}{F^{2}(x,y)\,g_{ij}(x,y)\,v^{i}v^{j}-[g_{ij}(x,y)\,y^{i}v^{j}]^{2}},$
where
$R^{i}_{\ k}=2\frac{\partial G^{i}}{\partial x^{k}}-y^{j}\frac{\partial^{2}G^{i}}{\partial x^{j}\partial y^{k}}+2G^{j}\frac{\partial^{2}G^{i}}{\partial y^{j}\partial y^{k}}-\frac{\partial G^{i}}{\partial y^{j}}\frac{\partial G^{j}}{\partial y^{k}}.$
If $F$ is a Riemannian metric, the flag curvature reduces to the sectional curvature.
The Ricci curvature of $y\in T_{x}M\backslash\{0\}$ is defined by
$\mathbf{Ric}(y):=\frac{1}{F^{2}(x,y)}\sum_{i}\mathbf{K}(y;e_{i}),$
where $\{e_{1},\ldots,e_{n}\}$ is a $g_{y}$-orthonormal basis for $T_{x}M$.
Given a locally Lipschitz curve $c:[0,1]\rightarrow M$, its length is defined by
$L_{F}(c)=\int_{0}^{1}F(c(t),c^{\prime}(t)){\operatorname{d}}t.$
The distance function $d_{F}:M\times M\rightarrow[0,\infty)$ is defined as $d_{F}(x_{1},x_{2}):=\inf L_{F}(c),$ where the infimum is taken over all locally Lipschitz curves $c:[0,1]\rightarrow M$ with $c(0)=x_{1}$ and $c(1)=x_{2}$. For any $x_{1},x_{2},x_{3}\in M$, one has that $d_{F}(x_{1},x_{2})\geq 0,$ with equality if and only if $x_{1}=x_{2}$, and $d_{F}(x_{1},x_{2})\leq d_{F}(x_{1},x_{3})+d_{F}(x_{3},x_{2})$. Usually $d_{F}(x_{1},x_{2})\neq d_{F}(x_{2},x_{1}),$ unless $F$ is reversible. In fact, one has for every $x_{1},x_{2}\in M$ that
$\frac{d_{F}(x_{1},x_{2})}{d_{F}(x_{2},x_{1})}\leq\lambda_{F}(M).$ (2.8)
Given $R>0$, the forward and backward metric balls $B^{+}_{x}(R)$ and $B^{-}_{x}(R)$ are defined by
$B^{+}_{x}(R):=\{z\in M\,:\,d_{F}(x,z)<R\},\quad B^{-}_{x}(R):=\{z\in M\,:\,d_{F}(z,x)<R\}.$
If $F$ is reversible, forward and backward metric balls coincide which are denoted by $B_{x}(R)$.
A Finsler manifold $(M,F)$ is called forward complete if every geodesic $t\mapsto\gamma(t)$, $t\in[0,1)$, can be extended to a geodesic defined on $t\in[0,+\infty)$; similarly, $(M,F)$ is called backward complete if every geodesic $t\mapsto\gamma(t)$, $t\in(0,1]$, can be extended to a geodesic on $t\in(-\infty,1]$. We say that $(M,F)$ is complete if it is both forward and backward complete. According to [2, 35], if $(M,F)$ is either forward or backward complete, then for every two points $x_{1},x_{2}\in M$, there exists a minimal geodesic $\gamma$ from $x_{1}$ to $x_{2}$ such that $d_{F}(x_{1},x_{2})=L_{F}(\gamma)$. Moreover, the closure of a forward (resp., backward) metric ball with finite radius is compact if $(M,F)$ is forward (resp., backward) complete.
The reverse Finsler metric $\overleftarrow{F}$ is defined as $\overleftarrow{F}(x,y):=F(x,-y)$. It is easy to check that $(M,F)$ is forward (resp., backward) complete if and only if $(M,\overleftarrow{F})$ is backward (resp., forward) complete. For this reason, we focus in the sequel only on forward complete Finsler manifolds.
2.2. Measure and polar coordinate system
In this subsection, let $(M,F,\operatorname{\mathfrak{m}})$ be a forward complete Finsler metric measure manifold ($\operatorname{\mathsf{FMMM}}$, for short), i.e., a forward complete Finsler manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$. In a local coordinate system ($x^{i}$), $\operatorname{d\mathfrak{m}}$ can be expressed as
$\operatorname{d\mathfrak{m}}=\sigma{\operatorname{d}}x^{1}\wedge\cdot\cdot\cdot\wedge{\operatorname{d}}x^{n},$ (2.9)
where $\sigma=\sigma(x)$ denotes the density function of $\operatorname{d\mathfrak{m}}$. In particular, the Busemann–Hausdorff measure $\operatorname{d\mathfrak{m}}_{BH}$ is defined by
$\operatorname{d\mathfrak{m}}_{BH}:=\frac{\operatorname{vol}(\mathbb{B}^{n})}{\operatorname{vol}(B_{x}M)}{\operatorname{d}}x^{1}\wedge\cdot\cdot\cdot\wedge{\operatorname{d}}x^{n}=\frac{\omega_{n}}{\operatorname{vol}(B_{x}M)}{\operatorname{d}}x^{1}\wedge\cdot\cdot\cdot\wedge{\operatorname{d}}x^{n},$
where $\mathbb{B}^{n}$ is the $n$-dimensional Euclidean unit ball, $\omega_{n}:=\operatorname{vol}(\mathbb{B}^{n})$ and $B_{x}M:=\{y\in T_{x}M\,:\,F(x,y)<1\}$.
The distortion $\tau$ and the S-curvature $\mathbf{S}$ of $(M,F,\operatorname{\mathfrak{m}})$ are defined by
$\tau(x,y):=\ln\frac{\sqrt{\det g_{ij}(x,y)}}{\sigma(x)},\qquad\mathbf{S}(x,y):=\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\tau(\gamma_{y}(t),{\gamma}^{\prime}_{y}(t)),\quad\forall\,y\in T_{x}M\setminus\{0\},$ (2.10)
respectively, where $t\mapsto\gamma_{y}(t)$ is a geodesic with $\gamma^{\prime}_{y}(0)=y$. According to Shen [27], we also have the explicit form
$\mathbf{S}(x,y)=\frac{\partial G^{m}}{\partial y^{m}}(x,y)-y^{k}\frac{\partial}{\partial x^{k}}(\ln\sigma(x)),$ (2.11)
where $G^{m}$ are the geodesic coefficients. Throughout this paper, we say that $\mathbf{S}\geq a$ for some constant $a\in\mathbb{R}$ if
$\mathbf{S}(x,y)\geq aF(x,y),\quad\forall\,(x,y)\in TM\backslash\{0\}.$
Let $(M,g)$ be a Riemannian manifold. Given an arbitrary smooth positive measure $\operatorname{\mathfrak{m}}$ on $M$, there always exists some $f\in C^{\infty}(M)$ such that $\operatorname{\mathfrak{m}}=e^{-f}\operatorname{vol}_{g}$. Thus, for any $x\in M$ and any $y\in T_{x}M\backslash\{0\}$,
$\tau(x,y)=f(x),\qquad\mathbf{S}(x,y)=g(y,\nabla f|_{x}),$ (2.12)
see Zhao [36]. According to (2.12), $S$-curvature in the weighted Riemannian setting cannot be of constant sign, unless it vanishes; similar situation occurs on a reversible Finsler manifold as well, see Zhao [36, (2.10)]. However, on irreversible Finsler manifolds, the $S$-curvature can be of constant sign, i.e., either positive or negative. Such phenomenon happens on every Funk metric space endowed with the Busemann–Hausdorff measure, when the $S$-curvature is a positive constant and it depends only on the dimension of the manifold, see Shen [26].
Given a point $o\in M$, set $S_{o}M:=\{y\in T_{o}M:F(o,y)=1\}$. Define the cut value $i_{y}$ of $y\in S_{o}M$ and the injectivity $\mathfrak{i}_{o}$ at $o$ as
$i_{y}:=\sup\{t>0:\text{the geodesic }\gamma_{y}|_{[0,t]}\text{ is globally minimizing}\},\quad\mathfrak{i}_{o}:=\inf_{y\in S_{o}M}i_{y}>0.$
In particular, if $(M,F)$ is a Cartan–Hadamard manifold (i.e., a simply connected, forward complete Finsler manifold with $\mathbf{K}\leq 0$), then $\mathfrak{i}_{x}=+\infty$ for every $x\in M$.
According to Zhao and Shen [37, Section 3], given a point $o\in M$, one can define the polar coordinate system $(r,y)$ around $o$, where $r(x):=d_{F}(o,x)$ and $y\in S_{o}M$. In particular, for any $y\in S_{o}M$ and $r\in(0,i_{y})$,
$\gamma_{y}(r)=(r,y),\quad\gamma^{\prime}_{y}(r)=\nabla r|_{(r,y)}.$ (2.13)
Moreover, it follows by Shen [27, Lemma 3.2.3] that the eikonal equations hold, i.e., $F(\nabla r)=F^{*}({\operatorname{d}}r)=1$ for $\operatorname{\mathfrak{m}}$-a.e. in $M$.
By the polar coordinate system $(r,y)$, the measure $\operatorname{\mathfrak{m}}$ can be disintegrated into the form
$\operatorname{d\mathfrak{m}}|_{(r,y)}=:\hat{\sigma}_{o}(r,y){\operatorname{d}}r\wedge{\operatorname{d}}\nu_{o}(y),$ (2.14)
where ${\operatorname{d}}\nu_{o}(y)$ is the Riemannian volume form of $S_{o}M$ induced by $g_{y}$. In particular, [37, Lemma 3.1] yields
$\lim_{r\rightarrow 0^{+}}\frac{\hat{\sigma}_{o}(r,y)}{r^{n-1}}=e^{-\tau(y)}.$ (2.15)
The following notations are useful in the sequel:
$\mathfrak{s}_{k}(t):=\left\{\begin{array}[]{lll}\frac{\sin(\sqrt{k}t)}{\sqrt{k}},&&\text{ if }k>0,\\ \ \ \ \ t,&&\text{ if }k=0,\\ \frac{\sinh(\sqrt{-k}t)}{\sqrt{-k}},&&\text{ if }k<0,\end{array}\right.\qquad\mathfrak{c}_{k}(t):=\frac{\operatorname{d}}{{\operatorname{d}}t}\mathfrak{s}_{k}(t)=\left\{\begin{array}[]{lll}{\cos(\sqrt{k}t)},&&\text{ if }k>0,\\ \ \ \ \ \ 1,&&\text{ if }k=0,\\ {\cosh(\sqrt{-k}t)},&&\text{ if }k<0.\end{array}\right.$
In view of (2.14), we recall a volume comparison result, see Zhao and Shen [37, Theorem 3.4].
Theorem 2.1.
(see [37]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}\geq(n-1)k$. For every $o\in M$, there holds
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{k}^{n-1}(r),\quad r\in(0,i_{y}),\quad\forall\,y\in S_{o}M,$ (2.16)
where $(r,y)$ is the polar coordinate system at $o$. In particular, we have
$\operatorname{\mathfrak{m}}[B^{+}_{o}(R)]\leq\int_{S_{o}M}\left(\int^{R}_{0}e^{-\tau({\gamma}_{y}^{\prime}(t))}\mathfrak{s}_{k}^{n-1}(r){\operatorname{d}}r\right){\operatorname{d}}\nu_{o}(y),$
for any $R>0$, with equality for some $R_{0}>0$ if and only if for any $y\in S_{o}M$,
$\mathbf{K}(\gamma_{y}^{\prime}(r);\cdot)\equiv k\quad\text{ for }0\leq r\leq R_{0}\leq\mathfrak{i}_{o},$
in which case
$\hat{\sigma}_{o}(r,y)=e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{k}^{n-1}(r),\quad r\in(0,R_{0}).$
Remark 2.1.
Due to (2.13), for $\operatorname{\mathfrak{m}}$-a.e. $x\in M$, the gradient $\nabla r|_{x}$ is the velocity of the minimal geodesic from $o$ to $x$ at the point $x$. By this observation and the same proof of [37, Theorem 3.4], it is not hard to check that the statements in Theorem 2.1 remain valid if the curvature condition is weakened to $\mathbf{Ric}(\nabla r)\geq(n-1)k.$ Moreover, if we additionally assume $\mathbf{S}(\nabla r)\geq(n-1)h$, then it follows by (2.13) that for any $y\in S_{o}M$,
$\tau({\gamma}_{y}^{\prime}(r))-\tau(y)=\tau({\gamma}_{y}^{\prime}(r))-\tau({\gamma}_{y}^{\prime}(0))=\int^{r}_{0}\mathbf{S}({\gamma}_{y}^{\prime}(s)){\operatorname{d}}s\geq(n-1)hr,$
which combined with (2.16) yields the key estimate in our forthcoming arguments, i.e.,
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{-k^{2}}^{n-1}(r)=e^{-\tau(y)-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r),\quad r\in(0,i_{y}).$ (2.17)
As in Huang, Kristály and Zhao [10], we recall the integral of distortion
$\mathscr{I}_{\operatorname{\mathfrak{m}}}(x):=\int_{S_{x}M}e^{-\tau(y)}{\operatorname{d}}\nu_{x}(y)<+\infty.$ (2.18)
This quantity expresses a close relation to the measure; in fact, $\mathscr{I}_{\operatorname{\mathfrak{m}}}$ is constant if and only if $\operatorname{\mathfrak{m}}=C\operatorname{\mathfrak{m}}_{BH}$ for some constant $C>0$. Moreover,
$\mathscr{I}_{\operatorname{\mathfrak{m}}_{BH}}(x)=\operatorname{vol}(\mathbb{S}^{n-1})=n\omega_{n},\quad\forall\,x\in M,$ (2.19)
where $\mathbb{S}^{n-1}$ is the $(n-1)$-dimensional Euclidean unit sphere.
Throughout this paper, $\chi_{U}$ denotes the characteristic function of a set $U\subset M$. As an application of Theorem 2.1 and Remark 2.1, we obtain the following result.
Lemma 2.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $o$ be a fixed point in $M$. Set $r(x):=d_{F}(o,x)$ and $\mu:=r^{-p}\operatorname{\mathfrak{m}}$ for some $p\in\mathbb{R}$. Thus,
  1. (i)
    if $p\in(-\infty,n)$, then $\int_{M}\chi_{\mathscr{K}}{\operatorname{d}}\mu<+\infty$ for any compact set $\mathscr{K}\subset M$;
  2. (ii)
    if $p\in[n,\infty)$, then $\int_{M}\chi_{B^{+}_{o}(R)}{\operatorname{d}}\mu=+\infty$ for any $R>0$.
Hence, $\mu=r^{-p}\operatorname{\mathfrak{m}}$ is a locally finite Borel measure if and only if $p\in(-\infty,n)$.
Proof.
(i) Given a compact set $\mathscr{K}\subset M$, since $r(x)$ is a continuous function, there is some $R\in(0,+\infty)$ such that $\mathscr{K}\subset{B^{+}_{o}(R)}$. Because the forward completeness implies the compactness of $\overline{B^{+}_{o}(R)}$, we may assume that
$\mathbf{Ric}|_{\overline{B^{+}_{o}(R)}}\geq-(n-1)k^{2},\quad\mathbf{S}|_{\overline{B^{+}_{o}(R)}}\geq(n-1)h,$
for some constants $k>0$ and $h\in\mathbb{R}$. It follows by Remark 2.1 that for any $y\in S_{o}M$,
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau({\gamma}_{y}^{\prime}(r))}\mathfrak{s}_{-k^{2}}^{n-1}(r)=e^{-\tau(y)-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r),\quad r\in(0,\min\{R,i_{y}\}).$ (2.20)
On the other hand, due to (2.15), there exists a small $\varepsilon_{0}\in(0,\min\{\mathfrak{i}_{o},R\})$ such that
$\hat{\sigma}_{o}(r,y)\leq 2e^{-\tau(y)}r^{n-1},\quad\forall\,r\in(0,\varepsilon_{0}),\ \forall\,y\in S_{o}M.$ (2.21)
By (2.14), (2.18), (2.20) and (2.21), we obtain
$\int_{M}\chi_{\mathscr{K}}{\operatorname{d}}\mu$ $\leq\int_{M}\chi_{B^{+}_{o}(R)}{\operatorname{d}}\mu\leq\int_{B^{+}_{o}(\varepsilon_{0})}r^{-p}\operatorname{d\mathfrak{m}}+\int_{B^{+}_{o}(R)\setminus B^{+}_{o}(\varepsilon_{0})}r^{-p}\operatorname{d\mathfrak{m}}$
$\leq 2\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int^{\varepsilon_{0}}_{0}r^{n-1-p}{\operatorname{d}}r\right)+\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int^{R}_{\varepsilon_{0}}e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{-p}{\operatorname{d}}r\right)$
$\leq\frac{2\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{n-p}\varepsilon_{0}^{n-p}+\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\frac{\varepsilon_{0}^{-p}}{(2k)^{n-1}}\int^{R}_{\varepsilon_{0}}e^{(n-1)(k-h)r}{\operatorname{d}}r<+\infty.$
(ii) For any $R>0$, the limit (2.15) furnishes a small $\delta_{0}\in(0,\min\{\mathfrak{i}_{o},R\})$ such that
$\hat{\sigma}_{o}(r,y)\geq\frac{1}{2}e^{-\tau(y)}r^{n-1},\quad\forall\,r\in(0,\delta_{0}),\ \forall\,y\in S_{o}M,$
which combined with (2.14) and (2.18) provides the estimate
$\displaystyle\int_{M}\chi_{B^{+}_{o}(R)}{\operatorname{d}}\mu$ $\geq\displaystyle\int_{B^{+}_{o}(\delta_{0})}r^{-p}\hat{\sigma}_{o}(r,y){\operatorname{d}}r{\operatorname{d}}\nu_{o}(y)\geq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{2}\int_{0}^{\delta_{0}}r^{n-1-p}{\operatorname{d}}r=+\infty,$
where we used the assumption $p\in[n,+\infty)$. ∎
3. Sobolev spaces over Finsler manifolds and approximation
Definition 3.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$. Given $u\in C^{\infty}_{0}(M)$ and $p\in[1,\infty)$, define a pseudo-norm as
$\|u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}:=\|u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}+\|{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}:=\left(\int_{M}|u|^{p}\operatorname{d\mathfrak{m}}\right)^{1/p}+\left(\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)^{1/p}.$ (3.1)
The Sobolev space $W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ is defined as the closure of $C^{\infty}_{0}(M)$ with respect to the backward topology induced by $\|\cdot\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}$, i.e.,
$u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})\quad\Longleftrightarrow\quad\exists\,(u_{n})\subset C^{\infty}_{0}(M)\text{ satisfying }\lim_{n\rightarrow\infty}\|u-u_{n}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}=0.$
Remark 3.1.
By the properties of $F^{*}$, it is not hard to verify that for any $u,v\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$,
  • $\|u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\geq 0$ with equality if and only if $u=0$;
  • $\|\alpha\,u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}=\alpha\|u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}$ for any $\alpha>0$;
  • $\|u+v\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\leq\|u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}+\|v\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}$.
Proposition 3.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$. If $W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ is not a vector space for some $p\in[1,\infty)$, then $\lambda_{F}(M)=+\infty$.
Proof.
Due to Remark 3.1, we clearly have that $\alpha_{1}u_{1}+\alpha_{2}u_{2}\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ for every $u_{1},u_{2}\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ and $\alpha_{1},\alpha_{2}\in\mathbb{R}_{+}$.
Since $W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ is not a vector space for some $p\in[1,\infty)$, there exists $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$ with $\|-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}=+\infty$, i.e., $\int_{M}F^{*p}(-{\operatorname{d}}u)\operatorname{d\mathfrak{m}}=+\infty$. Suppose by contradiction that $\lambda_{F}(M)<+\infty$. Thus (2.5) yields $\lambda_{F}(M)=\lambda_{F^{*}}(M)$ and hence,
$\int_{M}F^{*p}(-{\operatorname{d}}u)\operatorname{d\mathfrak{m}}\leq\lambda^{p}_{F}(M)\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}<+\infty,$
which is a contradiction. ∎
Proposition 3.2.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$. Suppose that there exist $p\in[1,\infty)$ and $q\in[1,p)$ satisfying
$\int_{M}\lambda^{\frac{qp}{p-q}}_{F}(x)\operatorname{d\mathfrak{m}}(x)<+\infty.$ (3.2)
Then $\operatorname{span}\{u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})\}$ is a linear subspace of $W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$.
Proof.
Since $\lambda_{F}(x)\geq 1$ for every $x\in M$, we have $\operatorname{\mathfrak{m}}(M)\leq\int_{M}\lambda^{\frac{qp}{p-q}}_{F}(x)\operatorname{d\mathfrak{m}}(x)<+\infty.$ Hence, without loss of generality, we may assume $\operatorname{\mathfrak{m}}(M)=1$ for convenience.
Given any $u\in W^{1,p}_{0}(M,\operatorname{\mathfrak{m}})$, there exists a sequence $(u_{k})\subset C^{\infty}_{0}(M)$ such that $\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$. Thus, Hölder’s inequality yields $\|u-u_{k}\|_{W^{1,q}_{\operatorname{\mathfrak{m}}}}\leq\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$, i.e., $u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. Moreover, (3.2) combined with Hölder’s inequality again gives
$\|{\operatorname{d}}(-u)-{\operatorname{d}}(-u_{k})\|_{L^{q}_{\operatorname{\mathfrak{m}}}}$ $=\left(\int_{M}F^{*q}({\operatorname{d}}u_{k}-{\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)^{1/q}\leq\left(\int_{M}\lambda_{F}^{q}F^{*q}({\operatorname{d}}u-{\operatorname{d}}u_{k})\operatorname{d\mathfrak{m}}\right)^{1/q}$
$\leq\left(\int_{M}\lambda_{F}^{\frac{qp}{p-q}}\operatorname{d\mathfrak{m}}\right)^{\frac{p-q}{qp}}\left(\int_{M}F^{*p}({\operatorname{d}}u-{\operatorname{d}}u_{k})\operatorname{d\mathfrak{m}}\right)^{\frac{1}{p}}\leq\|\lambda_{F}\|_{L^{\frac{qp}{p-q}}_{\operatorname{\mathfrak{m}}}}\|u-u_{k}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}$
$\rightarrow 0,$
which implies $-u\in W^{1,q}_{0}(M,\operatorname{\mathfrak{m}})$. ∎
In the rest of this section we consider the approximation by Lipschitz functions, which plays an important role in the proof of our main results.
Definition 3.2.
A function $f:M\rightarrow\mathbb{R}$ is called Lipschitz if there exists a constant $C\geq 0$ such that
$|f(x_{1})-f(x_{2})|\leq C\,d_{F}(x_{1},x_{2}),\quad\forall\,x_{1},x_{2}\in M.$
The function $f$ is also called $C$-Lipschitz. The minimal $C$ satisfying the above inequality is called the dilatation of $f$, denoted by $\mathsf{dil}(f)$.
For convenience, in the sequel we use $\mathsf{Lip}(M)$ to denote the collection of Lipschitz functions on $M$. Besides, let $\mathsf{Lip}_{0}(M):=C_{0}(M)\cap\mathsf{Lip}(M)$ be the collection of Lipschitz functions with compact support.
Lemma 3.1.
For a forward complete Finsler manifold $(M,F)$, one has $C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$.
Proof.
It suffices to show $C^{\infty}_{0}(M)\subset\mathsf{Lip}(M)$. In fact, for every $u\in C^{\infty}_{0}(M)$, given every $x_{1},x_{2}\in U:=\operatorname{supp}u$, the forward completeness of $(M,F)$ yields a unit speed minimal geodesic $\gamma(t)$, $t\in[0,d_{F}(x_{1},x_{2})]$ from $x_{1}$ to $x_{2}$. Since $U$ is compact, we have
$\lambda_{F}(U)<+\infty,\quad\max_{x\in U}F^{*}({\operatorname{d}}u|_{x})<+\infty.$
By letting $C:=\lambda_{F}(U)\max_{x\in U}F^{*}({\operatorname{d}}u|_{x})$ and using (2.4), we have
$|u(x_{1})-u(x_{2})|=$ $\left|\int^{d_{F}(x_{1},x_{2})}_{0}\frac{{\operatorname{d}}}{{\operatorname{d}}t}u(\gamma(t)){\operatorname{d}}t\right|=\left|\int^{d_{F}(x_{1},x_{2})}_{0}\langle\dot{\gamma}(t),{\operatorname{d}}u\rangle{\operatorname{d}}t\right|$
$\leq$ $\int^{d_{F}(x_{1},x_{2})}_{0}\left|\langle\dot{\gamma}(t),{\operatorname{d}}u\rangle\right|{\operatorname{d}}t\leq\lambda_{F}(U)\int^{d_{F}(x_{1},x_{2})}_{0}F(\dot{\gamma}(t))F^{*}({\operatorname{d}}u){\operatorname{d}}t$
$=$ $C\,d_{F}(x_{1},x_{2}),$
which implies $C^{\infty}_{0}(M)\subset\mathsf{Lip}(M)$. ∎
By Rademacher’s theorem and local coordinate systems, it follows that every Lipschitz function on $(M,F,\operatorname{\mathfrak{m}})$ is differentiable $\operatorname{\mathfrak{m}}$-a.e. In addition, we have the following approximation result which is useful in our proofs.
Proposition 3.3.
Let $(M,F,\operatorname{\mathfrak{m}})$ be a forward complete $\operatorname{\mathsf{FMMM}}$. For every $u\in\mathsf{Lip}_{0}(M)$, there exists a sequence of smooth Lipschitz functions $u_{l}\in C^{\infty}_{0}(M)\cap\mathsf{Lip}_{0}(M)$ such that
$\operatorname{supp}u\cup\operatorname{supp}u_{l}\subset\mathscr{K},\quad|u_{l}|\leq C,\quad F^{*}({\operatorname{d}}u_{l})\leq\mathsf{dil}(u_{l})\leq C,\quad u_{l}\rightrightarrows u,\quad\|u_{l}-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0,$
where $\mathscr{K}\subset M$ is a compact set and $C>0$ is a constant. In particular, if $u$ is nonnegative, so are $u_{l}$’s.
Proof.
Although the proof is standard, we provide it for readers’ convenience. Let $\mathbb{B}^{n}$ denote the standard Euclidean unit open ball in $\mathbb{R}^{n}$. The proof is divided into two steps.
Step 1. We show that for every $C$-Lipschitz function $h\in C_{0}(\mathbb{B}^{n})$ and every $\epsilon>0$, there exists a $C$-Lipschitz function $h_{\epsilon}\in C^{\infty}_{0}(\mathbb{B}^{n})$ such that
$h_{\epsilon}\rightrightarrows h,\quad\|h_{\epsilon}-h\|_{W^{1,p}(\mathbb{B}^{n})}\rightarrow 0,\quad\text{as }\epsilon\rightarrow 0^{+},$
where $\|\cdot\|_{W^{1,p}(\mathbb{B}^{n})}$ is the standard Sobolev norm on $\mathbb{B}^{n}$, see (3.1). In fact, let $h_{\epsilon}$ be the convolution of $h$ with a standard mollifier $\phi_{\epsilon}$, i.e., $h_{\epsilon}:=h*\phi_{\epsilon}$. Note that
$|h_{\epsilon}(x)-h_{\epsilon}(y)|\leq|h(\cdot)-h(\cdot+y-x)|*\phi_{\epsilon}\leq C|x-y|,\quad\forall\,x,y\in C_{0}(\mathbb{B}^{n}),$
i.e., $h_{\epsilon}$ is $C$-Lipschitz. The other properties of $h_{\epsilon}$ follow from the standard theory of Sobolev spaces. In particular, $h_{\epsilon}$ is nonnegative whenever $h\geq 0$.
Step 2. Since $\operatorname{supp}u$ is compact, there exists a finite coordinate covering $\{(V_{i},\varphi_{i})\}_{i=1}^{N<\infty}$ of $\operatorname{supp}u$ such that $V_{i}\Subset M$ and $\varphi_{i}(V_{i})=\mathbb{B}^{n}$. We may also assume that $\mathscr{K}:=\cup_{i=1}^{N}\overline{V_{i}}$ is a compact subset in $M$. Let $\{\eta_{i}\}_{i=1}^{N}$ be a smooth partition of unity subordinated to $\{V_{i}\}_{i=1}^{N}$. An argument similar to Bao, Chern and Shen [2, (6.2.3)] yields a constant $L\geq 1$ such that for all $i\in\{1,\ldots,N\}$,
$L^{-1}\leq\frac{F^{*}(\varphi^{*}\xi)}{|\xi|_{i}}\leq L,\ \forall\xi\in T^{*}\mathbb{B}^{n}\backslash\{0\},$ (3.3)
$|\eta_{i}(x_{1})-\eta_{i}(x_{2})|\leq L\,d_{F}(x_{1},x_{2}),\ \forall\,x_{1},x_{2}\in M,\ \max_{x\in U}F^{*}({\operatorname{d}}\eta_{i})\leq\sqrt{L},\ \lambda_{F}(\mathscr{K})\leq\sqrt{L},$ (3.4)
and
$L^{-1}d_{F}(x_{1},x_{2})\leq|\varphi_{i}(x_{1})-\varphi_{i}(x_{2})|_{i}\leq L\,d_{F}(x_{1},x_{2}),$ (3.5)
for every $x_{1},x_{2}\in V_{i}$ and $L^{-1}{\operatorname{d}}x_{i}\leq\operatorname{d\mathfrak{m}}|_{V_{i}}\leq L\,{\operatorname{d}}x_{i},$ where $|\cdot|_{i}$ and ${\operatorname{d}}x_{i}$ denote the Euclidean norm and the Lebesgue measure on $\varphi_{i}(V_{i})=\mathbb{B}^{n}$, respectively.
According to Lemma 3.1 and (3.5)1 (i.e., the first inequality in (3.5)), it is easy to check that $h_{i}:=(\eta_{i}u)\circ\varphi_{i}^{-1}$ is a $C_{1}$-Lipschitz function in $\mathbb{B}^{n}$ with compact support, where
$C_{1}:=\max_{1\leq i\leq N}\mathsf{dil}(h_{i})<+\infty.$
Due to Step 1, for each $k$ and $\epsilon>0$, there exists a $C_{1}$-Lipschitz function $h_{i,\epsilon}\in C^{\infty}_{0}(\mathbb{B}^{n})$ such that
$h_{i,\epsilon}\rightrightarrows h_{i},\quad\|h_{i,\epsilon}-h_{i}\|_{W^{1,p}(\mathbb{B}^{n})}\rightarrow 0,\quad\text{as }\epsilon\rightarrow 0^{+}.$ (3.6)
In particular, by the uniform convergence we may assume $|h_{i,\epsilon}|\leq C_{1}$, by choosing a larger $C_{1}$ if necessary.
Consider $h_{i,\epsilon}\circ\varphi_{i}\in C^{\infty}_{0}(V_{i})$; due to (3.5)1, they are $C_{1}L$-Lipschitz functions. Now set
$u_{\epsilon}:=\sum_{i=1}^{N}h_{i,\epsilon}\circ\varphi_{i}\in C^{\infty}_{0}(M),$
which is a $NC_{1}L$-Lipschitz function with $|u_{\epsilon}|\leq NC_{1}$ and $\operatorname{supp}u_{\epsilon}\subset\cup_{i=1}^{N}V_{i}\subset\mathscr{K}$. Moreover, by a direct calculation one has
$\left|u_{\epsilon}-u\right|$ $=\left|\sum_{i=1}^{N}h_{i,\epsilon}\circ\varphi_{i}-\sum_{i=1}^{N}h_{i}\circ\varphi_{i}\right|\leq\sum_{i=1}^{N}\left|h_{i,\epsilon}\circ\varphi_{i}-h_{i}\circ\varphi_{i}\right|,$
which together with (3.6) and (3.5)2 furnishes $u_{\epsilon}\rightrightarrows u$ and $\|u_{\epsilon}-u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0$. In addition, (3.3) yields
$F^{*}\left({\operatorname{d}}u_{\epsilon}-{\operatorname{d}}u\right)\leq\sum_{i=1}^{N}F^{*}\left({\operatorname{d}}(h_{i,\epsilon}\circ\varphi_{i})-{\operatorname{d}}(h_{i}\circ\varphi_{i})\right)\leq L\sum_{i=1}^{N}\left|{\operatorname{d}}h_{i,\epsilon}-{\operatorname{d}}h_{i}\right|_{i},$
which combined with (3.6)2 and (3.5)2 implies
$\|u_{\epsilon}-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0.$
In particular, if $u\geq 0$, then $h_{i}\geq 0$ and hence, $h_{i,\epsilon}\geq 0$ due to Step 1. Therefore, $u_{\epsilon}\geq 0$. We are done by letting $\epsilon=1/l$ and $C:=NC_{1}L$. ∎
4. Failure of functional inequalities: proof of Theorem 1.1
In this section we prove Theorem 1.1, considering in separate subsections the items (i)-(iv). Before doing this, we present an auxiliary result which will be useful in several steps.
Lemma 4.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $h\geq k\geq 0$ be constants. Suppose that there exists a point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1)h,$
where $r(x):=d_{F}(o,x)$. Let $u_{\alpha}(x):=-e^{-\alpha r(x)}$ for $\alpha>0$. Then the following statements hold::::
  1. (i)
    $\|u_{\alpha}\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}<+\infty$ for any $p\in[1,\infty);$
  2. (ii)
    $\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}<+\infty$ for any $p\in[1,n);$
  3. (iii)
    if $k>0$ then for any $p\in[1,\infty),$
    $\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}\leq\alpha^{-1}\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{pk^{n-1}},\quad\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}\leq\alpha^{p-1}\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{pk^{n-1}},$ (4.1)
    where $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)$ is defined by (2.18).
Proof.
Let $(r,y)$ be the polar coordinate system around $o\in M$. According to (2.17), we have
$\hat{\sigma}_{o}(r,y)\leq e^{-\tau(y)-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r),\quad\forall\,r\in(0,i_{y}),\ \forall\,y\in S_{o}M.$ (4.2)
The proof is divided into two cases.
Case 1. Suppose $k=0$, in which case $\mathfrak{s}_{-k^{2}}(r)=r$ and $h\geq 0$. Thus (4.2) combined with (2.14) yields
$\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}\leq$ $\int_{S_{o}M}\left(\int^{i_{y}}_{0}e^{-\tau(y)}e^{-p\alpha r-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right){\operatorname{d}}\nu_{o}(y)$
$\leq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$=$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr}r^{n-1}{\operatorname{d}}r$ (4.3)
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-p\alpha r}r^{n-1}{\operatorname{d}}r$ (4.4)
$=$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\frac{(n-1)!}{(p\alpha)^{n}}<+\infty.$
In addition, by the eikonal equation $F^{*}({\operatorname{d}}r)=1$ we have ${F}^{*}({\operatorname{d}}u_{\alpha})=\alpha e^{-\alpha r}=\alpha|u_{\alpha}|$, which implies
$\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}=\alpha^{p}\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}\leq\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\frac{(n-1)!}{p^{n}\alpha^{n-p}}<+\infty.$ (4.5)
Hence, (i) follows. Furthermore, when $p\in[1,n)$, a similar argument yields
$\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}=$ $\int_{M}e^{-p\alpha r}r^{-p}\operatorname{d\mathfrak{m}}\leq\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr}r^{n-1-p}{\operatorname{d}}r$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-p\alpha r}r^{n-1-p}{\operatorname{d}}r=\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\frac{\Gamma(n-p)}{(p\alpha)^{n-p}}<+\infty,$
which gives (ii).
Case 2. Suppose $k>0$, in which case $\mathfrak{s}_{-k^{2}}(r)=\frac{\sinh kt}{k}$. A computation analogous to (4.3) together with $h\geq k$ furnishes
$\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}\leq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$\leq$ $\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{k^{n-1}}\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr+(n-1)kr}{\operatorname{d}}r\leq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{k^{n-1}}\int^{+\infty}_{0}e^{-p\alpha r}{\operatorname{d}}r=\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{p\alpha k^{n-1}},$
which also yields
$\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}=\alpha^{p}\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}\leq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{pk^{n-1}}\alpha^{p-1}.$
Therefore, both (i) and (iii) hold. To prove (ii), we need the following estimates of $\hat{\sigma}_{o}(r,y)$. On the one hand, (4.2) yields
$\hat{\sigma}_{o}(r,y)\leq\frac{e^{-\tau(y)-(n-1)(h-k)r}}{(2k)^{n-1}},\quad\forall\,r\in(0,i_{y}),\ \forall\,y\in S_{o}M.$
On the other hand, (2.15) guarantees a small $\varepsilon_{0}\in(0,\mathfrak{i}_{o})$ such that
$\hat{\sigma}_{o}(r,y)\leq 2e^{-\tau(y)}r^{n-1},\quad\forall\,r\in(0,\varepsilon_{0}),\ \forall\,y\in S_{o}M.$
Thus, for any $p\in[1,n)$, an argument similar to (4.3) yields
$\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}$ $=\int_{B^{+}_{o}(\varepsilon_{0})}e^{-p\alpha r}r^{-p}\operatorname{d\mathfrak{m}}+\int_{M\setminus B^{+}_{o}(\varepsilon_{0})}e^{-p\alpha r}r^{-p}\operatorname{d\mathfrak{m}}$
$\leq 2\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{\varepsilon_{0}}_{0}r^{n-1-p}{\operatorname{d}}r+\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{(2k)^{n-1}}\int^{+\infty}_{\varepsilon_{0}}e^{-p\alpha r-(n-1)(h-k)r}r^{-p}{\operatorname{d}}r$
$\leq\frac{2\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{n-p}\varepsilon_{0}^{n-p}+\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{(2k)^{n-1}}\varepsilon_{0}^{-p}\int^{+\infty}_{\varepsilon_{0}}e^{-p\alpha r}{\operatorname{d}}r<+\infty,$
which concludes the proof. ∎
Remark 4.1.
Under the assumption of Lemma 4.1, one can prove in a standard manner that $u_{\alpha}\in W_{0}^{1,p}(M,\operatorname{\mathfrak{m}})$; since this fact is not used explicitly, we omit its proof.
4.1. Vanishing first eigenvalue of $p$-Laplacian
In this subsection we will study the first eigenvalue of the $p$-Laplacian in the Finsler setting. According to Shen [27], the divergence of a smooth vector field $X$ with respect to $\operatorname{\mathfrak{m}}$ is defined as
${\text{div}}_{\operatorname{\mathfrak{m}}}(X)\operatorname{d\mathfrak{m}}:={\operatorname{d}}(X\rfloor\operatorname{d\mathfrak{m}}).$
Given $p\in(1,\infty)$ and $u\in C^{\infty}(M)$, the $p$-Laplacian of $u$ (with respect to $\operatorname{\mathfrak{m}}$) is defined by
$\Delta_{p,\operatorname{\mathfrak{m}}}u:={\text{div}}_{\operatorname{\mathfrak{m}}}\left(F^{p-2}(\nabla u)\nabla u\right)$ (4.6)
on $\mathcal{U}:=\{x\in M\,:\,{\operatorname{d}}{u}|_{x}\neq 0\}$. Moreover, if $u\in W^{1,p}_{\operatorname{loc}}(M)$, the $p$-Laplacian of $u$ in the weak sense (with respect to $\operatorname{\mathfrak{m}}$) is defined by
$\int_{M}v\Delta_{p,\operatorname{\mathfrak{m}}}u\operatorname{d\mathfrak{m}}=-\int_{M}F^{p-2}(\nabla u)\langle\nabla u,{\operatorname{d}}v\rangle\operatorname{d\mathfrak{m}},$ (4.7)
for any $v\in C^{\infty}_{0}(M)$. In particular, the Laplacian $\Delta_{\operatorname{\mathfrak{m}}}u$ is exactly $\Delta_{2,\operatorname{\mathfrak{m}}}u$.
Let $p\in[1,\infty)$ and $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$; for further use, let us consider
$\Lambda_{p}(\cdot):=\frac{\displaystyle\int_{M}{F}^{*}({\operatorname{d}}\,\cdot\,)^{p}\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}|\cdot|^{p}\operatorname{d\mathfrak{m}}}.$
Lemma 4.2.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $h\geq k\geq 0$ be constants. Assume that there exists some point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1)h,$
where $r(x):=d_{F}(o,x)$. Then for any $p\in[1,\infty)$, one has
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\Lambda_{p}(u_{\alpha}),$ (4.8)
where $u_{\alpha}:=-e^{-\alpha r}$ for $\alpha>0$.
Proof.
We divide the proof into three steps.
Step 1. In this step, we are going to show that
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
Since $C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$, it is obvious that
$\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)\leq\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
For the reverse inequality, it suffices to show that for any $u\in\mathsf{Lip}_{0}(M)\backslash\{0\}$, there exists a sequence $u_{l}\in C^{\infty}_{0}(M)\backslash\{0\}$ such that
$\lim_{l\rightarrow\infty}\Lambda_{p}(u_{l})=\Lambda_{p}(u).$ (4.9)
Given $u\in\mathsf{Lip}_{0}(M)$, Proposition 3.3 provides a sequence of functions $(u_{l})\subset C^{\infty}_{0}(M)\cap\mathsf{Lip}_{0}(M)$ satisfying
$\operatorname{supp}u\cup\operatorname{supp}u_{l}\subset\mathscr{K},\quad u_{l}\rightrightarrows u,\quad\|u_{l}-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0,$ (4.10)
where $\mathscr{K}$ is some compact set in $M$. The uniform convergence in (4.10) implies
$\lim_{l\rightarrow\infty}\int_{M}|u_{l}|^{p}\operatorname{d\mathfrak{m}}=\int_{M}\lim_{l\rightarrow\infty}|u_{l}|^{p}\operatorname{d\mathfrak{m}}=\int_{M}|u|^{p}\operatorname{d\mathfrak{m}}.$ (4.11)
The $W^{1,p}$-convergence in (4.10) combined with (3.1) yields
$\|{\operatorname{d}}u_{l}-{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}\leq\|u_{l}-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0,\quad\text{ as }l\rightarrow\infty.$ (4.12)
As $\operatorname{supp}u\cup\operatorname{supp}u_{l}\subset\mathscr{K}$ and $\lambda_{F^{*}}(\mathscr{K})<+\infty$, we have
$\lim_{l\rightarrow\infty}\|{\operatorname{d}}u-{\operatorname{d}}u_{l}\|_{L^{p}_{\operatorname{\mathfrak{m}}}}\leq\lambda_{F^{*}}(\mathscr{K})\lim_{n\rightarrow\infty}\|{\operatorname{d}}u_{l}-{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}=0,$ (4.13)
Now the triangle inequality for $F^{*}$ implies that
$\|{\operatorname{d}}u_{l}\|_{L^{p}_{\operatorname{\mathfrak{m}}}}\leq\|{\operatorname{d}}u_{l}-{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}+\|{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}},\quad\|{\operatorname{d}}u\|_{L^{p}_{\operatorname{\mathfrak{m}}}}\leq\|{\operatorname{d}}u-{\operatorname{d}}u_{l}\|_{L^{p}_{\operatorname{\mathfrak{m}}}}+\|{\operatorname{d}}u_{l}\|_{L^{p}_{\operatorname{\mathfrak{m}}}},$
which combined with (4.12) and (4.13) yields
$\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}=\|{\operatorname{d}}u\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{n\rightarrow\infty}\|{\operatorname{d}}u_{l}\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{l\rightarrow\infty}\int_{M}F^{*p}({\operatorname{d}}u_{l})\operatorname{d\mathfrak{m}}.$ (4.14)
Thus, (4.9) follows directly by (4.11) and (4.14).
Step 2. In this step, we prove that for every $u_{\alpha}:=-e^{-\alpha r}$ with $\alpha>0$, there exists a sequence $(v_{\alpha,l})\subset\mathsf{Lip}_{0}(M)$ such that
$\Lambda_{p}(u_{\alpha})=\lim_{l\rightarrow\infty}\Lambda_{p}(v_{\alpha,l}).$ (4.15)
If $\sup_{x\in M}r(x)<+\infty$, then $M$ is compact and $u_{\alpha}\in C^{\infty}(M)=C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$. Hence, (4.15) is trivial by choosing $v_{\alpha,l}=u_{\alpha}$. In the sequel, we consider the case when $\sup_{x\in M}r(x)=+\infty$, i.e., $M$ is noncompact. Thus, for every small $\varepsilon\in(0,\min\{1,\mathfrak{i}_{o}\})$ and $\delta\in(0,e^{-\alpha})$, set
$u_{\alpha,\varepsilon}(x):=\begin{cases}-e^{-\alpha\varepsilon},&\text{ if }r(x)<\varepsilon;\\ \\ -e^{-\alpha r(x)},&\text{ if }r(x)\geq\varepsilon,\end{cases}\qquad u_{\alpha,\varepsilon,\delta}:=\min\{0,u_{\alpha,\varepsilon}+\delta\}.$
On the one hand, we claim that $u_{\alpha,\varepsilon,\delta}\in\mathsf{Lip}_{0}(M)$. In fact, it is easy to check that
$\operatorname{supp}u_{\alpha,\varepsilon,\delta}=\overline{\{u_{\alpha,\varepsilon,\delta}<0\}}=\overline{B^{+}_{o}\left(-\frac{\ln\delta}{\alpha}\right)},$
which is a compact set. In addition, since the construction implies
$F^{*}(\pm{\operatorname{d}}u_{\alpha,\varepsilon,\delta})\leq\max_{x\in\overline{B^{+}_{o}(-\alpha^{-1}\ln\delta)}}\max\{F^{*}(\pm{\operatorname{d}}u_{\alpha}|_{x})\}=:C<+\infty,$
an argument similar to the proof of Lemma 3.1 yields
$\left|u_{\alpha,\varepsilon,\delta}(x_{1})-u_{\alpha,\varepsilon,\delta}(x_{2})\right|\leq Cd_{F}(x_{1},x_{2}),\quad\forall\,x_{1},x_{2}\in M,$
which proves the claim.
On the other hand, set $\tilde{u}_{\alpha,\varepsilon}:=u_{\alpha,\varepsilon,\varepsilon}\in\mathsf{Lip}_{0}(M)$ for every $\varepsilon\ll 1$ and $v_{\alpha,l}:=\tilde{u}_{\alpha,1/l}\in\mathsf{Lip}_{0}(M)$. For (4.15), we only need to show
$\Lambda_{p}(u_{\alpha})=\lim_{\varepsilon\rightarrow 0^{+}}\Lambda_{p}(\tilde{u}_{\alpha,\varepsilon}).$ (4.16)
The above construction implies the following pointwise convergence
[…]
as $\varepsilon\to 0^{+}.$ Moreover, Lemma 4.1/(i) implies that $|u_{\alpha}|,F^{*}({\operatorname{d}}u_{\alpha})\in L^{p}(M,\operatorname{\mathfrak{m}})$, which together with the dominated convergence theorem and (4.17) yields
$\lim_{\varepsilon\rightarrow 0^{+}}\int_{M}|\tilde{u}_{\alpha,\varepsilon}|^{p}\operatorname{d\mathfrak{m}}=\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}},\quad\lim_{\varepsilon\rightarrow 0^{+}}\int_{M}F^{*p}({\operatorname{d}}\tilde{u}_{\alpha,\varepsilon})\operatorname{d\mathfrak{m}}=\int_{M}F^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}};$ (4.18)
thus, the limit in (4.16) follows.
Step 3. We now conclude the proof. For any $\alpha>0$ and $l\in\mathbb{N}$, Steps 1 and 2 yield
$\inf_{u\in C^{\infty}_{0}(M)}\Lambda_{p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)}\Lambda_{p}(u)\leq\Lambda(v_{\alpha,l}).$
By letting first $l\rightarrow\infty$ and then $\alpha\rightarrow 0^{+}$, and using (4.15), we get
$\inf_{u\in C^{\infty}_{0}(M)}\Lambda_{p}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\lim_{l\rightarrow\infty}\Lambda_{p}(v_{\alpha,l})=\liminf_{\alpha\rightarrow 0^{+}}\Lambda_{p}(u_{\alpha}),$
which is exactly (4.8). ∎
In view of (4.6) and (4.7), the first eigenvalue of the $p$($>1$)-Laplace operator $\Delta_{p,\operatorname{\mathfrak{m}}}$ is defined as
$\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}|u|^{p}\operatorname{d\mathfrak{m}}}=\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$ (4.19)
We have the following result, which proves a slightly more general version as Theorem 1.1/(i):
Theorem 4.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $h\geq k\geq 0$ be constants. If there exists some point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1)h,$
where $r(x):=d_{F}(o,x)$, then $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=0$ for all $p\in(1,\infty)$.
Proof.
For every $\alpha>0$, let $u_{\alpha}:=-e^{-\alpha r}$. Since $F^{*}({\operatorname{d}}r)=1$, we have
$\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}=\alpha^{p}\displaystyle\int_{M}e^{-p\alpha r}\operatorname{d\mathfrak{m}},\quad\displaystyle\int_{M}|u_{\alpha}|^{p}(x)\operatorname{d\mathfrak{m}}=\displaystyle\int_{M}e^{-p\alpha r}\operatorname{d\mathfrak{m}}.$ (4.20)
Then (4.19) combined with Lemma 4.2 and (4.20) yields
$\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)$ $=\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\Lambda_{p}(u_{\alpha})=\liminf_{\alpha\rightarrow 0^{+}}\alpha^{p}=0,$
which concludes the proof. ∎
Corollary 4.1.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1)h,$
for some constants $h\geq k\geq 0$. Thus $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=0$ for all $p\in(1,\infty)$.
Remark 4.2.
(i) Consider a noncompact Riemannian manifold equipped with the canonical Riemannian measure satisfying $\mathbf{Ric}\geq 0$. Corollary 4.1 implies the first eigenvalue of the Laplacian vanishes, which has been observed first by Cheng and Yau [6, Proposition 9].
(ii) Theorem 4.1 and Corollary 4.1 have been proved by Kristály [17, Theorem 1.3] in the special case when $(M,F,\operatorname{\mathfrak{m}})$ is the Funk metric on the unit Euclidean ball $\mathbb{B}^{n}$ endowed with the Busemann–Hausdorff measure.
The following result is a direct consequence of Corollary 4.1.
Corollary 4.2.
Let $(M,F)$ be an $n$-dimensional Cartan–Hadamard manifold with constant flag curvature $-k^{2}<0$. For each $p\in(1,\infty)$, the following statements hold::::
  1. (a)
    If there exists a positive smooth measure $\operatorname{\mathfrak{m}}$ such that $\mathbf{S}\geq(n-1)|k|$, then $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=0$.
  2. (b)
    If there exists a positive smooth measure $\operatorname{\mathfrak{m}}$ such that $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)>0$, then $\inf_{y\in SM}\mathbf{S}(y)<(n-1)|k|$.
Remark 4.3.
Let $(\mathbb{H}^{n},g)$ be the $n$-dimensional hyperbolic space with constant sectional curvature ${\bf K}=-k^{2}<0$, endowed with the canonical Riemannian measure $\operatorname{\mathfrak{m}}_{\rm can}$; one has the optimal spectral gap
$\lambda_{1,p,\operatorname{\mathfrak{m}}_{\rm can}}(\mathbb{H}^{n})=\frac{(n-1)^{2p}k^{2p}(p-1)^{p}}{p^{2p}}>0;$
see McKean [20] for $p=2$, and Farkas, Kajántó and Kristály [8] for generic $p>1$. Clearly, $\mathbf{S}=0$, which is consistent with Corollary 4.2/(b).
4.2. Failure of Hardy inequality
In this subsection, we investigate the Hardy inequality in the Finsler framework. As before, let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $o\in M$ be a fixed point. Given $p\in(1,\infty)$, set
$\mathcal{J}_{o,p}(u):=\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}},$ (4.21)
where $r(x):=d_{F}(o,x)$. In order to establish the $L^{p}$-Hardy inequality, it is equivalent to find
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
Due to Lemma 2.1/(ii), provided that $p\in[n,\infty)$, for every $u\in C^{\infty}_{0}(M)$ with $u(x)=c\neq 0$ for every $x\in B^{+}_{o}(r)$ and for some $r>0$, we have that
$\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}<+\infty,\quad\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}=+\infty,$ (4.22)
which implies
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=0,$
i.e., the Hardy inequality fails. Therefore, it is enough to consider the case when $p\in(1,n)$.
Lemma 4.3.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $h\geq k\geq 0$ be constants. Assume there exists a point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1)h,$
where $r(x):=d_{F}(o,x)$. Then for any $p\in(1,n)$ one has
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\mathcal{J}_{o,p}(u_{\alpha}),$ (4.23)
where $u_{\alpha}:=-e^{-\alpha r}$ for $\alpha>0,$ and $\mathcal{J}_{o,p}$ is defined by (4.21).
Proof.
The proof being similar to Lemma 4.2, we only sketch it, dividing into three steps.
Step 1. We show that
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
In view of $C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$, it suffices to show that for any $u\in\mathsf{Lip}_{0}(M)$, there exists a sequence $u_{l}\in C^{\infty}_{0}(M)\backslash\{0\}$ such that
$\lim_{l\rightarrow\infty}\mathcal{J}_{o,p}(u_{l})=\mathcal{J}_{o,p}(u).$ (4.24)
Given any $u\in\mathsf{Lip}_{0}(M)$, Proposition 3.3 guarantees the existence of a sequence $(u_{l})\subset C^{\infty}_{0}(M)\cap\mathsf{Lip}_{0}(M)$ satisfying
$\operatorname{supp}u\cup\operatorname{supp}u_{l}\subset\mathscr{K},\quad u_{l}\rightrightarrows u,\quad\|u_{l}-u\|_{W^{1,p}_{\operatorname{\mathfrak{m}}}}\rightarrow 0,$ (4.25)
where $\mathscr{K}\subset M$ is a compact set. The uniform convergence in (4.25) combined with Lemma 2.1/(i) yields
$\lim_{l\rightarrow\infty}\int_{M}\frac{|u_{l}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}=\lim_{l\rightarrow\infty}\int_{\mathscr{K}}|u_{l}|^{p}{\operatorname{d}}\mu=\int_{\mathscr{K}}\lim_{l\rightarrow\infty}|u_{l}|^{p}{\operatorname{d}}\mu=\int_{\mathscr{K}}|u|^{p}{\operatorname{d}}\mu=\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}},$
where ${\operatorname{d}}\mu=r^{-p}\operatorname{d\mathfrak{m}}$. Thus (4.24) follows because (4.14) remains valid.
Step 2. We are going to show that for every $u_{\alpha}:=-e^{-\alpha r}$ with $\alpha>0$, there exists a sequence $(v_{\alpha,l})\subset\mathsf{Lip}_{0}(M)$ such that
$\mathcal{J}_{o,p}(u_{\alpha})=\lim_{l\rightarrow\infty}\mathcal{J}_{o,p}(v_{\alpha,l}).$ (4.26)
It is enough to discuss the case $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and (4.26) is obvious by choosing $v_{\alpha,l}=u_{\alpha}\in C^{\infty}_{0}(M)$. Thus, for every small $\varepsilon>0$, set
$u_{\alpha,\varepsilon}:=\begin{cases}-e^{-\alpha\varepsilon},&\text{ if }r<\varepsilon,\\ \\ -e^{-\alpha r},&\text{ if }r\geq\varepsilon,\end{cases}\qquad\tilde{u}_{\alpha,\varepsilon}:=\min\{0,u_{\alpha,\varepsilon}+\varepsilon\},\qquad v_{\alpha,l}:=\tilde{u}_{\alpha,1/l}.$
Therefore, $\tilde{u}_{\alpha,\varepsilon}\in\mathsf{Lip}_{0}(M)$ follows by the same argument as in Step 2 of the proof of Lemma 4.2. Moreover, the dominated convergence theorem combined with (4.17) and Lemma 4.1/(i)&(ii) yields
$\lim_{\varepsilon\rightarrow 0^{+}}\int_{M}\frac{|\tilde{u}_{\alpha,\varepsilon}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}=\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}},\quad\lim_{\varepsilon\rightarrow 0^{+}}\int_{M}F^{*p}({\operatorname{d}}\tilde{u}_{\alpha,\varepsilon})\operatorname{d\mathfrak{m}}=\int_{M}F^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}},$
which implies $\lim_{\epsilon\rightarrow 0^{+}}\mathcal{J}_{o,p}(\tilde{u}_{\alpha,\varepsilon})=\mathcal{J}_{o,p}(u_{\alpha})$, concluding the proof of (4.24).
Step 3. By Steps 1 and 2, we obtain for every $l\in\mathbb{N}$ that
$\inf_{u\in C^{\infty}_{0}(M)}\mathcal{J}_{o,p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)}\mathcal{J}_{o,p}(u)\leq\mathcal{J}_{o,p}(v_{\alpha,l}),$
which together with (4.26) furnishes
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\lim_{l\rightarrow\infty}\mathcal{J}_{o,p}(v_{\alpha,l})=\liminf_{\alpha\rightarrow 0^{+}}\mathcal{J}_{o,p}(u_{\alpha}),$
completing the proof. ∎
Theorem 4.2.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $k,h$ be two constants with $h\geq k\geq 0$ and $h>0$. Suppose that there exists a point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1){h},$
where $r(x):=d_{F}(o,x)$. Thus, the $L^{p}$-Hardy inequality fails for all $p\in(1,n)$, i.e.,
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0.$ (4.27)
Proof.
Given $\alpha\in(0,1]$, set $u_{\alpha}(x):=-e^{-\alpha r(x)}$. Let $(r,y)$ denote the polar coordinate system around $o$. Due to (2.15), there exists an $\varepsilon\in(0,\mathfrak{i}_{o})$ such that
$\hat{\sigma}_{o}(r,y)\geq\frac{1}{2}e^{-\tau(y)}r^{n-1},\quad\forall\,r\in(0,\varepsilon),\ \forall\,y\in S_{o}M.$ (4.28)
Therefore, a direct calculation yields
$\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}$ $\geq\int_{S_{o}M}\int^{\varepsilon}_{0}\frac{e^{-\alpha pr}}{r^{p}}\frac{1}{2}e^{-\tau(y)}r^{n-1}{\operatorname{d}}\nu_{o}(y){\operatorname{d}}r\geq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)e^{-\alpha p\varepsilon}}{2}\int^{\varepsilon}_{0}{r^{n-p-1}}{\operatorname{d}}r$
$=\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{2e^{\alpha p\varepsilon}(n-p)}\varepsilon^{n-p}.$ (4.29)
The rest of the proof is divided into two cases.
Case 1: $k>0$. Thus, (4.21) together with (4.23), (4.29) and (4.1) yields
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}$
$\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\frac{2e^{\alpha p\varepsilon}(n-p)\alpha^{p-1}}{pk^{n-1}\varepsilon^{n-p}}$
$=$ $0,$
which is exactly (4.27).
Case 2: $k=0$. Due to ${F}^{*}({\operatorname{d}}u_{\alpha})=\alpha e^{-\alpha r}=\alpha|u_{\alpha}|$, an argument similar to (2.17) yields
$\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}=$ $\alpha^{p}\int_{M}|u_{\alpha}|^{p}\operatorname{d\mathfrak{m}}$
$\leq$ $\alpha^{p}\int_{S_{o}M}\left(\int^{+\infty}_{0}e^{-\tau(y)}e^{-p\alpha r-(n-1)hr}r^{n-1}{\operatorname{d}}r\right){\operatorname{d}}\nu_{o}(y)$
$=$ $\alpha^{p}\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-p\alpha r-(n-1)hr}r^{n-1}{\operatorname{d}}r$
$=$ $\alpha^{p}\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\frac{(n-1)!}{(\alpha p+(n-1)h)^{n}}<+\infty.$ (4.30)
Thus, (4.23) together with (4.29) and (4.30) implies
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u_{\alpha}|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}$
$\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\frac{2e^{\alpha p\varepsilon}(n-p)(n-1)!\,\alpha^{p}}{(\alpha p+(n-1)h)^{n}\varepsilon^{n-p}}$
$=$ $0,$
which concludes the proof. ∎
A direct consequence of Theorem 4.2 reads as follows.
Corollary 4.3.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h}$
where $k,h$ are two constants with $h\geq k\geq 0$ and $h>0$. Then for every point $o\in M$ and every $p\in(1,n)$, the $L^{p}$-Hardy inequality fails, i.e., (4.27) holds.
Remark 4.4.
(i) A closer inspection of the proof of Theorem 4.2 clearly shows that if $h=0$, the last limit would be +$\infty$+ ∞ since $p<n$. This is not surprising since in the Riemannian setting, the $L^{p}$-Hardy inequality is valid on every $n$-dimensional Cartan–Hadamard manifold equipped with the canonical Riemannian measure, due to the vanishing $S$-curvature (2.12), see Yang, Su and Kong [34] and Kristály [16]. On the other hand, there are several Finsler manifolds satisfying the assumptions of Corollary 4.3; see Section 6 for details.
(ii) In Corollary 4.3, the noncompactness of $M$ is unnecessary. Indeed, if $M$ is compact, then the constant functions belong to $C^{\infty}_{0}(M)=C^{\infty}(M)$, which immediately furnishes (4.27).
4.3. Failure of Heisenberg–Pauli–Weyl uncertainty principle
In this subsection, we study the Heisenberg–Pauli–Weyl principle.
Theorem 4.3.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $k,h$ be two constants with $h>k\geq 0$. Assume that there exists a point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1){h},$
where $r(x):=d_{F}(o,x)$. Then the Heisenberg–Pauli–Weyl uncertainty principle fails, i.e.,
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*}({\operatorname{d}}u)^{2}\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}u^{2}r^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$ (4.31)
Proof.
We only sketch the proof as it is similar to that of Theorem 4.2. For simplicity of notations, set
$\mathcal{H}_{o}(u):=\frac{\left(\displaystyle\int_{M}F^{*}({\operatorname{d}}u)^{2}\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}u^{2}r^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}.$ (4.32)
By the same method as employed in the proof Lemma 4.3, one can show that
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{H}_{o}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\mathcal{H}_{o}(u_{\alpha}),$ (4.33)
where $u_{\alpha}:=-e^{-\alpha r}$ for $\alpha>0$. Thus it suffices to show $\lim_{\alpha\rightarrow 0^{+}}\mathcal{H}_{o}(u_{\alpha})=0$.
Since $h>k\geq 0$, it is easy to check that
$\lim_{r\rightarrow+\infty}e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{2}=0,$
which yields a $\delta=\delta(n,k,h)>0$ such that
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{2}<1\quad\text{for }r\in(\delta,\infty).$ (4.34)
Relation (2.17) combined with (4.34) yields
$\int_{M}u_{\alpha}^{2}r^{2}\operatorname{d\mathfrak{m}}\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int_{0}^{\delta}+\int_{\delta}^{+\infty}e^{-2\alpha r-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{2}{\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\mathfrak{s}_{-k^{2}}^{n-1}(\delta)\int^{\delta}_{0}r^{2}{\operatorname{d}}r+\int_{\delta}^{+\infty}e^{-2\alpha r}{\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\frac{\delta^{3}\mathfrak{s}_{-k^{2}}^{n-1}(\delta)}{3}+\frac{1}{2\alpha}\right).$ (4.35)
On the other hand, since (4.28) is always valid, we obtain
$\int_{M}u_{\alpha}^{2}\operatorname{d\mathfrak{m}}$ $\geq\int_{S_{o}M}\int^{\varepsilon}_{0}e^{-2\alpha r}\frac{1}{2}e^{-\tau(y)}r^{n-1}{\operatorname{d}}\nu_{o}(y){\operatorname{d}}r\geq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{2}\int^{\varepsilon}_{0}e^{-2\alpha r}r^{n-1}{\operatorname{d}}r\geq\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{2ne^{2\alpha\varepsilon}}\varepsilon^{n},$
which together with (4.35) and (4.32) furnishes
$\lim_{\alpha\rightarrow 0^{+}}\mathcal{H}_{o}(u_{\alpha})=\lim_{\alpha\rightarrow 0^{+}}\frac{\alpha^{2}\displaystyle\int_{M}u_{\alpha}^{2}r^{2}\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}u_{\alpha}^{2}\operatorname{d\mathfrak{m}}}\leq\lim_{\alpha\rightarrow 0^{+}}2ne^{2\alpha\varepsilon}\varepsilon^{-n}\alpha^{2}\left(\frac{\delta^{3}\mathfrak{s}_{-k^{2}}^{n-1}(\delta)}{3}+\frac{1}{2\alpha}\right)=0,$
completing the proof. ∎
Corollary 4.4.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}\geq-(n-1)k^{2}$ and $\mathbf{S}\geq(n-1){h},$ where $k,h$ are two constants satisfying $h>k\geq 0$. Then for every point $o\in M$, the Heisenberg–Pauli–Weyl uncertainty principle fails, i.e., (4.31) holds.
We now discuss the indispensability of our assumption $h>k\geq 0$ in Theorem 4.3 and Corollary 4.4.
Remark 4.5.
(i) One cannot relax the assumption $h>k\geq 0$ to $h=0$ (and $k=0$). Indeed, according to Huang, Kristály and Zhao [10, Theorem 1.2], if a Cartan–Hadamard manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$ satisfies
$\mathbf{S}\leq 0,\quad\lambda_{F}(M)<+\infty,$ (4.36)
then the following Heisenberg–Pauli–Weyl uncertainty principle holds:
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{H}_{o}(u)\geq\frac{1}{\lambda_{F}(M)^{2}}\frac{n^{2}}{4}.$
Note that there are large classes of Cartan–Hadamard manifolds satisfying (4.36); for example, every Minkowski space endowed with the Busemann–Hausdorff measure. Therefore, the positivity condition $h>0$ in Theorem 4.3 cannot be relaxed, which is in a perfect concordance also with the setting of [10, Theorem 1.3]. In fact, as we shall see in Section 5, the positivity of the $S$-curvature leads to the infinity of the reversibility.
(ii) One cannot relax the assumption $h>k\geq 0$ to $h=k>0$. To see this, let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional noncompact forward complete $\operatorname{\mathsf{FMMM}}$. Assume that there is $o\in M$ with
$\mathbf{K}(\nabla r)\leq-k^{2},\quad\mathbf{S}(\nabla r)\leq(n-1)k,$
for some $k>0$, where $r(x):=d_{F}(o,x)$. Then we have the (non-sharp) Heisenberg–Pauli–Weyl uncertainty principle:
$\left(\int_{M}\max\{F^{*2}(\pm{\operatorname{d}}u)\}\operatorname{d\mathfrak{m}}\right)\left(\int_{M}r^{2}u^{2}\operatorname{d\mathfrak{m}}\right)\geq\frac{1}{4}\left(\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2},\quad\forall\,u\in C^{\infty}_{0}(M).$ (4.37)
Indeed, due to Wu and Xin [33, Lemma 3.3 & Theorem 4.1], we have $\operatorname{\mathfrak{m}}$-a.e. in $M$ that
$\Delta r\geq(n-1)k{\coth(kr)}-(n-1)k=\frac{2(n-1)k}{e^{2kr}-1}>0.$
The divergence theorem, relation (2.4), the eikonal equation and Hölder’s inequality yield
$\left(2\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}$ $\leq\left(2\int_{M}(1+r\Delta_{\operatorname{\mathfrak{m}}}r)u^{2}\operatorname{d\mathfrak{m}}\right)^{2}=\left(\int_{M}u^{2}\Delta_{\operatorname{\mathfrak{m}}}(r^{2})\operatorname{d\mathfrak{m}}\right)^{2}$
$=16\left(\int_{M}ur\langle\nabla r,{\operatorname{d}}u\rangle\operatorname{d\mathfrak{m}}\right)^{2}\leq 16\left(\int_{M}|u|r\max\{F^{*}(\pm{\operatorname{d}}u)\}\operatorname{d\mathfrak{m}}\right)^{2}$
$\leq 16\left(\int_{M}|u|^{2}r^{2}\operatorname{d\mathfrak{m}}\right)\left(\int_{M}\max\{F^{*2}(\pm{\operatorname{d}}u)\}\operatorname{d\mathfrak{m}}\right),$
which gives (4.37).
Remark 4.6.
The interpretation of the classical Heisenberg–Pauli–Weyl uncertainty principle in quantum mechanics is that the position and momentum of a given particle cannot be accurately determined simultaneously; this phenomenon can be captured in the presence of the ’gap’ $n^{2}/4$ in relation (1.3) (or the value $1/4$ in the non-sharp inequality (4.37)). However, in our geometric context provided by Theorem 4.3, the theoretical possibility of simultaneous detection of position and momentum of a particle still persist due to the gap-vanishing fact.
4.4. Failure of Caffarelli–Kohn–Nirenberg interpolation inequality
Theorem 4.4.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ and let $k,h$ be nonnegative constants with $h>k\geq 0$. Suppose that there exists a point $o\in M$ such that
$\mathbf{Ric}(\nabla r)\geq-(n-1)k^{2},\quad\mathbf{S}(\nabla r)\geq(n-1){h},$
where $r(x):=d_{F}(o,x)$. Then the Caffarelli–Kohn–Nirenberg interpolation inequality fails, i.e.,
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*2}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0,$ (4.38)
for any $p,q\in\mathbb{R}$ with
$0<q<2<p<+\infty,\quad 2<n<\frac{2(p-q)}{p-2}.$
Proof.
For every $\alpha>0$, let us consider the function
$u_{\alpha}(x):=-\left[1+(\alpha\,r(x))^{2-q}\right]^{\frac{1}{2-p}}.$
We split the proof into three steps.
Step 1. We first show that
$\int_{M}F^{*2}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}<+\infty,\quad\int_{M}\frac{|u_{\alpha}|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}<+\infty,\quad\int_{M}\frac{|u_{\alpha}|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}<+\infty.$ (4.39)
Since $\lim_{r\rightarrow 0^{+}}\mathfrak{s}^{n-1}_{-k^{2}}(r)/r^{n-1}=1$, there exists an $\varepsilon=\varepsilon(n,k)\in(0,\mathfrak{i}_{o})$ such that
$r^{2-2q}\mathfrak{s}^{n-1}_{-k^{2}}(r)\leq 2r^{n-2q+1},\quad\forall r\in(0,\varepsilon).$ (4.40)
On the other hand, since $\eta:=h-k>0$, a similar argument yields $\delta=\delta(n,q,k,h)\in(\varepsilon,\infty)$ with the property that
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{(n-1)}(r)r^{2-2q}\leq e^{-\frac{(n-1)\eta}{2}r},\quad\forall\,r\in(\delta,\infty).$ (4.41)
An argument similar to (2.17) combined with (4.40) and (4.41) implies that
$\int_{M}\frac{|u_{\alpha}|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}=$ $\int_{M}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{2p-2}{2-p}}r^{2-2q}\operatorname{d\mathfrak{m}}$
$\leq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int_{0}^{i_{y}}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{2p-2}{2-p}}r^{2-2q}e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\int^{+\infty}_{0}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r$
$=$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int^{\varepsilon}_{0}+\int_{\varepsilon}^{\delta}+\int_{\delta}^{\infty}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(2\int^{\varepsilon}_{0}r^{n-2q+1}{\operatorname{d}}r+\int_{\varepsilon}^{\delta}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r+\int_{\delta}^{\infty}e^{-\frac{(n-1)\eta}{2}r}{\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\frac{2\varepsilon^{n-2q+2}}{n-2q+2}+\int_{\varepsilon}^{\delta}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r+\frac{2e^{-\frac{(n-1)\eta\delta}{2}}}{(n-1)\eta}\right)$
=:absent:=:= : $C_{1}(n,q,k,h,\delta,\varepsilon)<+\infty.$ (4.42)
By an analogous computation one can also show $\int_{M}\frac{|u_{\alpha}|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}<+\infty$. Moreover, a direct calculation yields
${\operatorname{d}}u_{\alpha}=\alpha^{2-q}\left(\frac{2-q}{p-2}\right)\left[1+(\alpha\,r)^{2-q}\right]^{\frac{p-1}{2-p}}r^{1-q}{\operatorname{d}}r,$
which together with (4.42) implies
$\int_{M}F^{*2}({\operatorname{d}}u_{\alpha})\operatorname{d\mathfrak{m}}=\alpha^{2(2-q)}\left(\frac{2-q}{p-2}\right)^{2}\int_{M}\frac{|u_{\alpha}|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}<+\infty.$ (4.43)
This completes the proof of (4.39).
Step 2. We show that there exists a sequence $u_{\alpha,l}\in\mathsf{Lip}_{0}(M)$ such that
$\mathcal{G}_{o,p,q}(u_{\alpha})=\lim_{l\rightarrow\infty}\mathcal{G}_{o,p,q}(u_{\alpha,l}),$ (4.44)
where
$\mathcal{G}_{o,p,q}(u):=\frac{\left(\displaystyle\int_{M}F^{*2}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}.$ (4.45)
As before, it is enough to consider the case when $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and the sequence would be $u_{\alpha,n}:=u_{\alpha}\in C^{\infty}(M)=C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$. Thus, for small $\varepsilon\in(0,1)$ set
$u_{\alpha,\varepsilon}:=\begin{cases}-[1+(\alpha\,\varepsilon)^{2-q}]^{\frac{1}{2-p}},&\text{if }r<\varepsilon,\\ \\ -[1+(\alpha\,r)^{2-q}]^{\frac{1}{2-p}},&\text{if }r\geq\varepsilon,\end{cases}\qquad v_{\alpha,\varepsilon}:=\min\{u_{\alpha,\varepsilon}+\varepsilon,0\},\qquad u_{\alpha,l}:=v_{\alpha,1/l},\ l\in\mathbb{N}.$
On can easily show that $v_{\alpha,\varepsilon},u_{\alpha,l}\in\mathsf{Lip}_{0}(M)$ since $u_{\alpha}\in C^{\infty}(M)$ with $\lim_{r(x)\rightarrow\infty}u_{\alpha}(x)=0$. In addition, the construction shows the pointwise convergences
$|u_{\alpha,l}|\nearrow|u_{\alpha}|,\quad F^{*}({\operatorname{d}}u_{\alpha,l})\nearrow F^{*}({\operatorname{d}}u_{\alpha})\ \ {\rm as}\ \ l\rightarrow\infty.$
Thus (4.44) can be derived by these limits, (4.39) and the dominated convergence theorem.
Step 3. On the one hand, by the same method as employed in the proof of Lemma 4.3 one can show
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{G}_{o,p,q}(u)\leq\liminf_{\alpha\rightarrow 0^{+}}\mathcal{G}_{o,p,q}(u_{\alpha}).$ (4.46)
On the other hand, since (4.28) remains valid, we have
$\int_{M}\frac{|u_{\alpha}|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}=$ $\int_{M}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{p}{2-p}}r^{-q}\operatorname{d\mathfrak{m}}$
$\geq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int^{\varepsilon}_{0}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{p}{2-p}}r^{-q}\frac{1}{2}r^{n-1}{\operatorname{d}}r\right)$
$\geq$ $\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)}{2}\left[1+(\alpha\varepsilon)^{2-q}\right]^{\frac{p}{2-p}}\int^{\varepsilon}_{0}r^{n-1-q}{\operatorname{d}}r$
$=$ $\frac{\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\varepsilon^{n-q}}{2(n-q)}\left[1+(\alpha\,\varepsilon)^{2-q}\right]^{\frac{p}{2-p}}=:C_{2}({\alpha},n,p,q,k,h,\varepsilon)>0.$ (4.47)
Now it follows from (4.46), (4.42), (4.43) and (4.47) that
$\inf_{u\in C^{\infty}_{0}(M)}\mathcal{G}_{o,p,q}(u)\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\mathcal{G}_{o,p,q}(u_{\alpha})$
$=$ $\liminf_{\alpha\rightarrow 0^{+}}\alpha^{2(2-q)}\left(\frac{2-q}{p-2}\right)^{2}\left[\frac{\displaystyle\int_{M}\frac{|u_{\alpha}|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u_{\alpha}|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}}\right]^{2}$
$=$ $\liminf_{\alpha\rightarrow 0^{+}}\alpha^{2(2-q)}\left(\frac{2-q}{p-2}\right)^{2}\left(\frac{C_{1}}{C_{2}}\right)^{2}=0,$
i.e., (4.38) follows. ∎
Remark 4.7.
We note that if $h=k>0$ and $q\in(1,2)$, then (4.41) and $\int_{M}\frac{|u_{\alpha}|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}<+\infty$ still hold and hence, the Caffarelli–Kohn–Nirenberg interpolation inequality fails. However, it remains open the question when $h=k>0$ and $q\in(0,1)$, as we cannot guarantee the validity of (4.41).
Corollary 4.5.
Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}\geq-(n-1)k^{2}$ and $\mathbf{S}\geq(n-1){h},$ where $k,h$ are two constants satisfying $h>k\geq 0$. Then for every point $o\in M$, the Caffarelli–Kohn–Nirenberg interpolation inequality fails, i.e., (4.38) holds for any $p,q\in\mathbb{R}$ with
$0<q<2<p<+\infty,\quad 2<n<\frac{2(p-q)}{p-2}.$
Proof of Theorem 1.1.
The statements are direct consequences of Corollaries 4.1, 4.3, 4.4 and 4.5, respectively. ∎
5. Influence of $S$-curvature on geometric aspects: proof of Theorem 1.2
5.1. Infinite reversibility
In this subsection we prove Theorem 1.2, which establish a surprising relationship between the $S$-curvature and reversibility. To do this, we recall the following Hardy inequality.
Theorem 5.1.
(Zhao [36, Theorem 3.8]) Let $(M,F,\operatorname{\mathfrak{m}})$ be an $n$-dimensional forward complete noncompact $\operatorname{\mathsf{FMMM}}$ with
$\mathbf{K}\leq 0,\quad\mathbf{S}\geq 0.$
[…]
In particular, the constant $\left(\frac{n+\beta}{p}\right)^{p}$ is sharp if $\lambda_{F}(M)=1$.
Proof of Theorem 1.2.
[…]
For any $p\in(1,n)$, Theorem 5.1 (by choosing $\beta=-p$) combined with (5.1) furnishes
$\lambda^{p}_{F}(M)\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\geq\left(\frac{n-p}{p\lambda_{F}(M)}\right)^{p}\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}},\quad\forall\,u\in C^{\infty}_{0}(M),$
which implies
$\inf_{u\in C^{\infty}_{0}(M)\setminus\{0\}}\frac{\displaystyle\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}\geq\left(\frac{n-p}{p\lambda^{2}_{F}(M)}\right)^{p}>0.$
However, according to Corollary 4.3, the left hand side of the above estimate is zero, which is a contradiction. Therefore, we necessarily have $\lambda_{F}(M)=+\infty$, which concludes the proof. ∎
Corollary 5.1.
Let $(M,F)$ be an $n$-dimensional Cartan–Hadamard manifold equipped with a smooth positive measure $\operatorname{\mathfrak{m}}$. Assume that $\mathbf{K}\geq-k^{2}$ and $\mathbf{S}\geq(n-1){h}$ for some constants $k,h$ with $h\geq k\geq 0$ and $h>0$. Then $\lambda_{F}(M)=+\infty$.
6. Examples & Counterexamples: Funk metric spaces
The purpose of this section is twofold. First, in subsection 6.1 we investigate Funk metric spaces, a large class of non-Riemannian Finsler manifolds that provide the geometric background for the applicability of our main results, Theorems 1.1-1.2. Second, in subsection 6.2, as we already noticed in the Introduction, we show that the combined assumption on the Ricci and $S$-curvatures cannot be replaced by the weighted Ricci curvature ${\bf Ric}_{N}$, in spite of the fact that the latter notion is a combination of $\mathbf{Ric}$ and $\mathbf{S}$.
Throughout this section, let $\Omega\subset\mathbb{R}^{n}$ be a bounded strongly convex domain with $\mathbf{0}\in\Omega$, let $(x^{i})$ denote the standard coordinate system of $\mathbb{R}^{n}$ and let $\operatorname{vol}(\cdot)$ denote the Euclidean volume.
6.1. Funk metric spaces: applicability of Theorems 1.1-1.2
Definition 6.1.
A Finsler metric $F$ on $\Omega$ is called a (general) Funk metric if
$x+\frac{y}{F(x,y)}\in\partial\Omega.$
for every $x\in\Omega$ and $y\in T_{x}\Omega\backslash\{0\}$. In the sequel, the pair $(\Omega,F)$ is called a (general) Funk metric space.
The main result of this subsection can be formulated as follows:
Theorem 6.1.
Any $n$-dimensional Funk metric space $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$ endowed with the Busemann–Hausdorff measure verifies the assumptions of Theorems 1.1-1.2. In particular, the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality fail on $(\Omega,F,\operatorname{\mathfrak{m}}_{BH})$, respectively.
Before giving the proof of Theorem 6.1, we recall/collect some useful properties of Funk metric spaces. We first notice that every Funk metric is projectively flat. In addition, according to Li [19] and Shen [31], a Finsler metric $F$ on $\Omega$ is a Funk metric if and only if there is a unique Minkowski norm $\phi=\phi(y)$ on $\mathbb{R}^{n}$ such that
$\partial\Omega=\phi^{-1}(1),\quad F(x,y)=\phi(y+xF(x,y)).$ (6.1)
In particular,
$\Omega=\{x\in\mathbb{R}^{n}\,:\,\phi(x)<1\}.$ (6.2)
Having this characterization of Funk metrics, we state the following set of properties; since we did not find them in the literature in this general form, we provide also their proof.
Proposition 6.1.
Let $(\Omega,F)$ be a Funk metric space and let $\phi=\phi(y)$ the Minkowski norm satisfying (6.1). Then the following statements hold::::
  1. (i)
    the distance function is given by
    $d_{F}(x_{1},x_{2})=\ln\left[\frac{\phi(z-x_{1})}{\phi(z-x_{2})}\right],\quad\forall\,x_{1},x_{2}\in\Omega,$ (6.3)
    where $z$ is the intersection point of the ray $l_{x_{1}x_{2}}(t):=x_{1}+t(x_{2}-x_{1})$, $t\geq 0$, with $\partial\Omega$, i.e.,
    $z=x_{1}+\lambda(x_{2}-x_{1})\in\partial\Omega\quad\text{ for some }\lambda>1;$
  2. (ii)
    for any $x\in\Omega$, one has
    $d_{F}(\mathbf{0},x)=-\ln(1-\phi(x)),\qquad d_{F}(x,\mathbf{0})=\ln(1+\phi(-x));$ (6.4)
  3. (iii)
    $(\Omega,F)$ is forward complete but not backward complete with $\lambda_{F}(\Omega)=+\infty$. In particular,
    $\lambda_{F}(x)=\sup_{y\in T_{x}\Omega\setminus\{0\}}\frac{F(x,-y)}{F(x,y)}\in\left[\frac{1+\phi(x)}{1-\phi(x)},\frac{\lambda_{\phi}+\phi(x)}{1-\phi(x)}\right],\quad\forall\,x\in\Omega.$ (6.5)
Proof.
(i) Given $x_{1},x_{2}\in\Omega$, let $t\mapsto c(t)$ be a geodesic from $x_{1}$ to $x_{2}$ with $c(0)=x$ and $c^{\prime}(0)=y$. Due to the projective flatness of $(\Omega,F)$, for an appropriate parameter $t=t(s)$, we may assume that $c(s):=x+sy.$
Since $x+\frac{1}{F(x,y)}y\in\partial\Omega$ by Definition 6.1, we have $x+sy\notin\partial\Omega$ if $s>\frac{1}{F(x,y)}$, which implies $c(s)\in\Omega$ for $s\in[0,\frac{1}{F(x,y)}]$. Then along the direction $y$, the corresponding point in $\partial\Omega$ is given by
$\partial\Omega\ni z:=x+\frac{y}{F(x,y)}=c(s)+\frac{y}{F(c(s),\dot{c}(s))}=x+sy+\frac{y}{F(c(s),\dot{c}(s))},$
which implies
$F(c(s),\dot{c}(s))=\frac{F(x,y)}{1-F(x,y)s}=\frac{F(y)}{1-F(y)s},$
where $F(y)$ is $F(x,y)$ for short. Thus, formula (6.3) giving the distance ${d}_{F}(x_{1},x_{2})$ between $x_{1}=c(s_{1})$ to $x_{2}=c(s_{2})$ is
${d}_{F}(x_{1},x_{2})=$ $\int_{s_{1}}^{s_{2}}F(c(s),\dot{c}(s)){\operatorname{d}}s=\int_{s_{1}}^{s_{2}}\frac{F(y)}{1-F(y)s}{\operatorname{d}}s=\ln\left[\frac{1-F(y)s_{1}}{1-F(y)s_{2}}\right]=\ln\left[\frac{\frac{1}{F(y)}-s_{1}}{\frac{1}{F(y)}-s_{2}}\right]$
$=$ $\ln\left[\frac{(\frac{1}{F(y)}-s_{1})\phi(y)}{(\frac{1}{F(y)}-s_{2})\phi(y)}\right]=\ln\left[\frac{\phi(\frac{y}{F(y)}-s_{1}y)}{\phi(\frac{y}{F(y)}-s_{2}y)}\right]=\ln\left[\frac{\phi(x+\frac{y}{F(y)}-x-s_{1}y)}{\phi(x+\frac{y}{F(y)}-x-s_{2}y)}\right]$
$=$ $\ln\left[\frac{\phi(z-x_{1})}{\phi(z-x_{2})}\right].$
(ii) Given $x_{1},x_{2}\in\Omega$, let $z:=x_{1}+\lambda(x_{2}-x_{1})\in\partial\Omega$ for some $\lambda>1$. Then (6.1) and (6.3) yield
11 11 $=\phi(z)=\phi(x_{1}+\lambda(x_{2}-x_{1})),$ (6.6)
$d_{F}(x_{1},x_{2})$ $=\ln\left[\frac{\phi(\lambda(x_{2}-x_{1}))}{\phi\left((\lambda-1)(x_{2}-x_{1})\right)}\right]=\ln\frac{\lambda}{\lambda-1}.$ (6.7)
By choosing $x_{1}=\mathbf{0}$ and $x_{2}=x$, we derive $\lambda=1/\phi(x)$ from (6.6), which together with (6.7) yields
$d_{F}(\mathbf{0},x)=\ln\frac{\frac{1}{\phi(x)}}{\frac{1}{\phi(x)}-1}=-\ln(1-\phi(x)).$
On the other hand, when $x_{1}=x$ and $x_{2}=\mathbf{0}$, from (6.6) we have $\lambda=\frac{1}{\phi(-x)}+1$; thus (6.7) gives
$d_{F}(x,\mathbf{0})=\ln(1+\phi(-x)).$ (6.8)
(iii) Given $x_{1},x_{2}\in\Omega$, the triangle inequality combined with (6.4) yields
$d_{F}(x_{1},x_{2})\geq d_{F}(\mathbf{0},x_{2})-d_{F}(\mathbf{0},x_{1})=-\ln(1-\phi(x_{2}))+\ln(1-\phi(x_{1})).$ (6.9)
Let $t\mapsto c(t)$, $t\in[0,1],$ be the unit speed geodesic from $c(0)=x_{1}$ to $c(1)=x_{2}$ (i.e., the ray $l_{x_{1}x_{2}}$). We extend $c$ to the point $z=x_{1}+\lambda(x_{2}-x_{1})\in\partial\Omega$; hence,
$\lim_{c(t)\rightarrow z}\phi(c(t))=\phi(z)=1,$
which combined with (6.9) yields
$d_{F}(c(0),c(t))=d_{F}(x_{1},c(t))\geq-\ln(1-\phi(c(t)))-\ln(1+\phi(x_{1}))\rightarrow+\infty,$
as $\phi(c(t))\rightarrow 1^{-}$, which means that every unit speed geodesic can be forward extended to $t=+\infty$. Thus, $(\Omega,F)$ is forward complete. On the other hand, (6.8) implies $d_{F}(x,\mathbf{0})\rightarrow\ln 2$ as $\phi(-x)\rightarrow 1^{-}$. Hence, ($\Omega$, $F$) cannot be backward complete because the bounded closed backward ball $\overline{B^{-}_{\mathbf{0}}(\ln 2)}=\Omega$ is not compact.
Given $x\in\Omega$ and considering it as a (tangent) vector in $\mathbb{R}^{n}$, by (2.4) we have
$\langle x,\xi\rangle\leq\phi(x)\phi^{*}(\xi),\quad\forall\,\xi\in T^{*}_{x}\Omega=\mathbb{R}^{n},$ (6.10)
with equality if and only if $\xi=\alpha\mathfrak{L}_{\phi}(x)$ for some $\alpha\geq 0$; here, $\mathfrak{L}_{\phi}$ stands for the Legendre transformation induced by $\phi$. It costs no generality to assume that $\phi^{*}(\mathfrak{L}_{\phi}(x))\leq\phi^{*}(-\mathfrak{L}_{\phi}(x))$. According to [11, 28], the co-metric $F^{*}$ of $F$ is given by
$F^{*}(x,\xi)=\phi^{*}(\xi)-\langle x,\xi\rangle,$
where $\phi^{*}$ is the co-metric of $\phi$. Therefore, by (2.5) one has
$\lambda_{F}(x)=$ $\lambda_{F^{*}}(x)=\sup_{\xi\in T^{*}_{x}\Omega\setminus\{0\}}\frac{F^{*}(x,-\xi)}{F^{*}(x,\xi)}=\sup_{\xi\in T^{*}_{x}\Omega\setminus\{0\}}\frac{\phi^{*}(-\xi)+\langle x,\xi\rangle}{\phi^{*}(\xi)-\langle x,\xi\rangle}$
$\geq$ $\frac{\phi^{*}(-\mathfrak{L}_{\phi}(x))+\phi(x)\phi^{*}(\mathfrak{L}_{\phi}(x))}{\phi^{*}(\mathfrak{L}_{\phi}(x))-\phi(x)\phi^{*}(\mathfrak{L}_{\phi}(x))}=\frac{\frac{\phi^{*}(-\mathfrak{L}_{\phi}(x))}{\phi^{*}(\mathfrak{L}_{\phi}(x))}+\phi(x)}{1-\phi(x)}\geq\frac{1+\phi(x)}{1-\phi(x)}.$
In particular, $\lambda_{F}(x)\to+\infty$ as $\phi(x)\rightarrow 1^{-}$, thus $\lambda_{F}(\Omega)=+\infty$. Moreover,
$\lambda_{F}(x)\leq\sup_{\xi\in T^{*}_{x}\Omega\setminus\{0\}}\frac{\frac{\phi^{*}(-\xi)}{\phi^{*}(\xi)}+\phi(x)}{1-\phi(x)}=\frac{\sup_{\xi\in T^{*}_{x}\Omega\setminus\{0\}}\frac{\phi^{*}(-\xi)}{\phi^{*}(\xi)}+\phi(x)}{1-\phi(x)}=\frac{\lambda_{\phi}+\phi(x)}{1-\phi(x)},$
which finishes the proof. ∎
After this preparatory part, we have:
Proof of Theorem 6.1.
By Proposition 6.1/(iii), $(\Omega,F)$ is forward complete. In addition, according to Shen [26], one has that $(\Omega,F)$ is a Cartan–Hadamard manifold with
  1. (i)
    constant flag curvature $\mathbf{K}=-\frac{1}{4};$
  2. (ii)
    constant $S$-curvature $\mathbf{S}=\frac{n+1}{2};$
  3. (iii)
    $\operatorname{d\mathfrak{m}}_{BH}=\sigma{\operatorname{d}}x$, where
    $\sigma\equiv\frac{\operatorname{vol}(\mathbb{B}^{n})}{\operatorname{vol}(\Omega)}=\frac{\omega_{n}}{\operatorname{vol}(\Omega)}.$ (6.11)
In particular, we may choose
$k=\frac{1}{2}\ \ {\rm and}\ \ h=\frac{n+1}{2(n-1)}$
in the assumptions of Theorems 1.1 &\&& 1.2, which guarantee the validity of (1.5) and (1.7). ∎
Remark 6.1.
In the special case of Funk metrics, due to (6.4), the main test function in the proof of Theorem 1.1 reduces to
$u_{\alpha}(x)=-e^{-\alpha d_{F}(\mathbf{0},x)}=-(1-\phi(x))^{\alpha},\ \alpha>0.$
This explicit form of $u_{\alpha}$ as well as the density $\sigma$ in (6.11) of the measure $\operatorname{\mathfrak{m}}_{BH}$ on $(\Omega,F)$ allow a direct, alternative verification of the failure of functional inequalities in Theorem 1.1. Such approach was performed in Kristály [17] to prove the vanishing of the first eigenvalue of the Laplacian in the special case when $\Omega=\mathbb{B}^{n}$ and $\phi(x)=|x|$, in which case the Funk metric has the explicit form appearing in (1.6).
We conclude this subsection by recalling the interpolation metric considered in Kristály and Rudas [18]; namely, for $a\in[0,1]$, let $F_{a}:T\mathbb{B}^{n}\to[0,\infty)$ be given by
$F_{a}(x,y)=\frac{\sqrt{|y|^{2}-|x|^{2}|y|^{2}+\langle x,y\rangle^{2}}+a\langle x,y\rangle}{1-|x|^{2}},\quad x\in\mathbb{B}^{n},\ y\in T_{x}\mathbb{B}^{n}=\mathbb{R}^{n}.$ (6.12)
It is clear that $F_{a}$ interpolates between the Funk metric (1.6) for $a=1$ with ${\bf K}=-1/4$ and ${\bf S}_{a}=(n+1)/2,$ and the Hilbert/Klein metric for $a=0$ with ${\bf K}=-1$ and ${\bf S}_{a}=0$, where ${\bf S}_{a}$ denotes the $S$-curvature of the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$. Note that $(\mathbb{B}^{n},F_{a})$ is a Randers-type projectively flat Finsler space with the reversibility $\lambda_{F_{a}}(\mathbb{B}^{n})=\frac{1+a}{1-a}$ for $a\in[0,1)$ and $\lambda_{F_{1}}(\mathbb{B}^{n})=+\infty$. Moreover, in Kajántó and Kristály [14, Theorems 1.1&1.2] – up to an isometry – it is proved that the flag curvature and $S$-curvature of $F_{a}$ with $a\in(0,1)$, can be estimated as
[…]
moreover, these estimates are sharp for every $a\in(0,1)$. In particular, for every $a\in[0,1)$ one has that $\inf_{(x,y)\in T\mathbb{B}^{n}}{\bf S}_{a}(x,y)=0$, thus Theorem 1.2 does not apply (i.e., $0=h<k$). However, by using Theorem 6.1, elementary estimates as in [18] and the validity of famous functional inequalities on Riemannian Cartan–Hadamard manifolds, see [10, 16], we can state:
Proposition 6.2.
Let $a\in[0,1]$, and consider the interpolated Funk-type metric space $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ endowed with its natural Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{a}$. Then the Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg inequality hold on $(\mathbb{B}^{n},F_{a},\operatorname{\mathfrak{m}}_{a})$ – with possibly not sharp constants – if and only if $a\in[0,1)$.
6.2. Counterexamples under $\mathbf{Ric}_{N}\geq-K^{2}$
Our main results are valid under the separate curvature assumptions
$\mathbf{Ric}\geq-(n-1)k^{2},\quad\mathbf{S}\geq(n-1){h}.$ (6.13)
Recall that there is another kind of Ricci curvature on such a space, which is a combination of $\mathbf{Ric}$ and $\mathbf{S}$, see Ohta [21]. More precisely, for a $\operatorname{\mathsf{FMMM}}$ $(M,F,\mathfrak{m})$, the weighted Ricci curvature $\mathbf{Ric}_{N}$ is defined as follows: given $N\in[n,\infty]$, for any unit vector $y\in SM$,
$\mathbf{Ric}_{N}(y)=\left\{\begin{array}[]{lll}\mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\dot{\gamma}_{y}(t))-\frac{\mathbf{S}^{2}(y)}{N-n},&&\text{ for }N\in(n,\infty),\\ \\ \underset{L\downarrow n}{\lim}\mathbf{Ric}_{L}(y),&&\text{ for }N=n,\\ \\ \mathbf{Ric}(y)+\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\mathbf{S}(\dot{\gamma}_{y}(t)),&&\text{ for }N=\infty.\end{array}\right.$ (6.19)
Weighted Ricci curvature is the realization of $\operatorname{\mathsf{CD}}(K,N)$ condition in the Finsler framework, which is a Finslerian version of the Bakry–Émery Ricci tensor. In view of this notion, it is a natural question whether or not the separated condition (6.13) can be replaced by $\mathbf{Ric}_{N}\geq-K^{2}$ for some $K\geq 0$. In fact, the following examples show that the statements of Theorems 1.1 &\&& 1.2 fail under $\mathbf{Ric}_{N}\geq-K^{2}$.
Example 6.1.
(see Li [19]) There are infinitely many reversible Minkowski norms on Euclidean spaces. In the sequel, we provide two such classes:
  • (a)
    Given $m,n\in\mathbb{Z}^{+}$, then
    $\phi(y)=\sqrt[2m]{\sum^{n}_{i=1}(y^{i})^{2m}},\ y=(y_{1},...,y_{n})\in\mathbb{R}^{n},$
    is a reversible Minkowski norm on $\mathbb{R}^{n}$.
  • (b)
    Let $|y|$ and $|\tilde{y}|$ be the Euclidean norms on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$, respectively. Then
    $\phi(y,\tilde{y})=\sqrt{\sqrt{|y|^{4}+|\tilde{y}|^{4}}+|y|^{2}+|\tilde{y}|^{2}},\ (y,\tilde{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m},$
    is a reversible Minkowski norm on $\mathbb{R}^{n}\times\mathbb{R}^{m}$.
In general, given a reversible Minkowski norm $\phi$ on the Euclidean space $\mathbb{R}^{n}$, the triple $(\mathbb{R}^{n},\phi,\mathscr{L}^{n})$ is a complete noncompact $\operatorname{\mathsf{FMMM}}$ with $\mathbf{Ric}_{N}\equiv 0$ for all $n\in[N,\infty]$. However, according to Huang, Kristaly and Zhao [10], the $L^{p}$-Hardy inequality, Heisenberg–Pauli–Weyl uncertainty principle and Caffarelli–Kohn–Nirenberg interpolation inequality are valid as well as sharp with the same optimal constants as in the Euclidean case; we also notice that $\mathbf{S}\equiv 0$ for such spaces.
Example 6.2 (Hilbert/Klein metric).
Let $F=F(x,y)$ be a Funk metric on a bounded strongly convex $\Omega\subset\mathbb{R}^{n}$, see Definition 6.1. The corresponding Hilbert metric (or Klein metric) is given as
$\tilde{F}(x,y)=\frac{1}{2}\left\{F(x,y)+F(x,-y)\right\}.$ (6.20)
We consider the triplet $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$, where $\operatorname{\mathfrak{m}}_{BH}$ is the Busemann–Hausdorff measure of $F$ (not $\tilde{F}$). We first observe that $\tilde{F}$ is reversible, i.e., $\lambda_{\tilde{F}}(\Omega)=1$. Then, according to Shen [26], $(\Omega,\tilde{F},\operatorname{\mathfrak{m}}_{BH})$ is a complete, $n$-dimensional projectively flat $\operatorname{\mathsf{FMMM}}$ of constant flag curvature $\widetilde{\mathbf{K}}=-1$. In addition, due to Ohta [22], one has that
$\widetilde{\mathbf{Ric}}_{N}\geq-(n-1)-\frac{(n+1)^{2}}{N-n},\quad\forall\,N\in(n,\infty].$ (6.21)
These properties clearly show that in Theorem 1.2 we cannot replace the combined curvature conditions with the lower boundedness of the weighted Ricci curvature.
Remark 6.2.
We can provide a short, alternative proof to (6.21), by using our previous arguments combined with some well-known facts about Hilbert metrics.
First, we observe that $\widetilde{\mathbf{Ric}}=-(n-1)$, as $\widetilde{\mathbf{K}}=-1$. Moreover, one has that
$-(n+1)\leq\tilde{\mathbf{S}}\leq n+1.$ (6.22)
Indeed, according to Okada [23], one has that every Funk metric satisfies
$\frac{\partial F}{\partial x^{k}}=F\frac{\partial F}{\partial y^{k}}.$ (6.23)
Let $P=P(x,y)$ denote the projective factor of the Funk metric $F$. Thus, (6.23) yields
$P=\frac{y^{k}}{2F}\frac{\partial F}{\partial x^{k}}=\frac{y^{k}}{2}\frac{\partial F}{\partial y^{k}}=\frac{1}{2}F.$
Then the projective factor of the Hilbert metric $\tilde{F}$ is given by
$\tilde{P}(x,y)$ $=\frac{y^{k}\tilde{F}_{x^{k}}(x,y)}{2\tilde{F}(x,y)}=\frac{P(x,y)F(x,y)-P(x,-y)F(x,-y)}{F(x,y)+F(x,-y)}$
$=\frac{F^{2}(x,y)-F^{2}(x,-y)}{2\left[F(x,y)+F(x,-y)\right]}=\frac{1}{2}\left[F(x,y)-F(x,-y)\right],$
which implies that for every $y\neq 0$, we have
$\frac{\tilde{P}^{2}(x,y)}{\tilde{F}^{2}(x,y)}=\frac{[F(x,y)-F(x,-y)]^{2}}{[F(x,y)+F(x,-y)]^{2}}\in[0,1).$ (6.24)
Consider the Busemann–Hausdorff measure $\operatorname{\mathfrak{m}}_{BH}$ of $F$, whose density function $\sigma(x)$ is a constant, see (6.11). Since $\tilde{G}^{m}=\tilde{P}y^{m}$, relation (2.11) yields
$\tilde{\mathbf{S}}(x,y)$ $=\frac{\partial\tilde{G}^{m}(x,y)}{\partial y^{m}}-y^{m}\frac{\partial}{\partial x^{m}}(\ln\sigma(x))=y^{m}\frac{\partial\tilde{P}(x,y)}{\partial y^{m}}+n\tilde{P}(x,y)=({n+1})\tilde{P}(x,y),$ (6.25)
which combined with (6.24) yields (6.22).
Finally, according to Troyanov [32, p. 103], we have
$\frac{1}{\tilde{F}^{2}}\left(\tilde{P}^{2}-\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}\right)=\tilde{\mathbf{K}}=-1.$ (6.26)
Thus, for any $y\in\widetilde{S\Omega}:=\{v\in T\Omega\,:\,\tilde{F}(v)=1\}$, by (2.6), (6.25) and (6.26), we obtain
$\left.\frac{\operatorname{d}}{{\operatorname{d}}t}\right|_{t=0}\tilde{\mathbf{S}}(\dot{\gamma}_{y}(t))$ $=\frac{\partial\tilde{\mathbf{S}}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{\mathbf{S}}}{\partial y^{m}}\tilde{G}^{m}(y)=(n+1)\left(\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}-2\frac{\partial\tilde{{P}}}{\partial y^{m}}\tilde{G}^{m}(y)\right)$
$=(n+1)\left(\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}-2\tilde{P}\frac{\partial\tilde{{P}}}{\partial y^{m}}y^{m}\right)$
$=(n+1)\left(\frac{\partial\tilde{P}}{\partial x^{k}}y^{k}-2\tilde{P}^{2}\right)=(n+1)(\tilde{F}^{2}-\tilde{P}^{2}).$ (6.27)
Therefore, it follows by (6.19), (6.25) and (6.2) that for any unit vector $y\in\widetilde{S\Omega}$,
$\widetilde{\mathbf{Ric}}_{N}(y)=\left\{\begin{array}[]{lll}2-(n+1)\tilde{P}^{2}-\frac{(n+1)^{2}}{N-n}\tilde{P}^{2},&&\text{ for }N\in(n,\infty),\\ \\ 2-(n+1)\tilde{P}^{2},&&\text{ for }N=\infty,\end{array}\right.$
which combined with relation (6.24) implies the required estimate (6.21).